### v1

# GRADA: Graph-based Reranker against Adversarial Documents Attack

Abstract
Retrieval Augmented Generation (RAG) frameworks improve the accuracy of large language models (LLMs) by integrating external knowledge from retrieved documents, thereby overcoming the limitations of models’ static intrinsic knowledge. However, these systems are susceptible to adversarial attacks that manipulate the retrieval process by introducing documents that are adversarial yet semantically similar to the query. Notably, while these adversarial documents resemble the query, they exhibit weak similarity to benign documents in the retrieval set. Thus, we propose a simple yet effective Graph-based Reranking against Adversarial Document Attacks (GRADA) framework aiming at preserving retrieval quality while significantly reducing the success of adversaries. Our study evaluates the effectiveness of our approach through experiments conducted on five LLMs: GPT-3.5-Turbo, GPT-4o, Llama3.1-8b-Instruct, Llama3.1-70b-Instruct, and Qwen2.5-7b-Instruct. We use three datasets to assess performance, with results from the Natural Questions dataset demonstrating up to an 80% reduction in attack success rates while maintaining minimal loss in accuracy.
GRADA: Graph-based Reranker against Adversarial Documents Attack
Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3, Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu5 1University of Melbourne, 2University of Edinburgh, 3University College London 4University of Manchester, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@manchester.ac.uk, qiongkai.xu@mq.edu.au
1 Introduction
Large Language Models (LLMs; Brown et al., 2020) have demonstrated remarkable performance across a wide range of natural language processing tasks, including question answering Fourrier et al. (2024), text summarization Graff et al. (2003); Rush et al. (2015), and information retrieval Yates et al. (2021). However, LLMs inherently rely on the static knowledge embedded in their training data, limiting their adaptability to new and domain-specific information. Retrieval-Augmented Generation (RAG; Lewis et al., 2020) was introduced to bridge this gap by integrating external retrieval modules, allowing LLMs to access and incorporate relevant, up-to-date knowledge.
Figure 1: An example of adversarial RAG attack which exploits query-document similarity by prepending the poisonous document with the query.
While RAG enhances the flexibility of LLMs, it also introduces new vulnerabilities. Adversaries can exploit retrieval mechanisms by injecting manipulated documents into the corpus Zhong et al. (2023); Clop and Teglia (2024); Greshake et al. (2023); Pasquini et al. (2024), subtly altering rankings to mislead LLM outputs. As shown in Figure 1, these adversarial documents mimic query-relevant patterns, making them difficult to detect while degrading the reliability of retrieval-based LLM systems. In practice, LLMs are used in the search engine to produce direct answers to user’s query, called Answer Engine Optimization (AEO)  Yalçın and Köse (2024). Adversarial documents can be created using the previous methods to achieve higher rankings in the search results. Also, guide the search engine to produce malicious context that can potentially harm the user Hammond (2024); Venkit et al. (2024).
Figure 2: An overview of GRADA. A vanilla RAG pipeline concatenates all retrieved documents along with the question as the input to the LLM. However, the accuracy of this pipeline can be easily harmed by malicious passages. In contrast, GRADA uses a graph-based approach to isolate and filter out malicious passages before passing the retrieved documents as the LLM input.
Existing noise filtering methods, such as Hybrid List Aware Transformer Reranking (HLATR, Zhang et al., 2022) and BAAI General Embeddings (BGE-reranker, Xiao et al., 2023), focus on improving document relevance by filtering out generic noise or low-quality content. However, these methods are ineffective against adversarial attacks that exploit query-document similarity patterns to evade detection. In addition, a recent study has reviewed current graph-based reranking methods. It shows a potential path of using graphs in future information retrieval Zaoad et al. (2025) tasks but the effects on adversarial documents remains unknown. On the other hand, specialized adversarial defenses, such as keyword filtering and decoding aggregation Xiang et al. (2024), can successfully remove adversarial content but at the cost of discarding valuable benign documents, ultimately weakening retrieval performance. This trade-off highlights the need for a more nuanced defense mechanism that can distinguish between adversarial and benign documents without compromising retrieval quality.
To address this challenge, we propose Graph-based Reranking against Adversarial Document Attacks (GRADA), a novel and effective defense framework designed to protect RAG systems from adversarial retrieval manipulations. Our key insight is that adversarial documents, while optimized for high query similarity, exhibit weaker semantic coherence with genuinely relevant documents in the retrieval set. Leveraging this property, we construct a graph where each retrieved document is represented as a node, and edges capture document-document similarity relationships. By propagating ranking scores through this graph structure, our approach prioritizes clusters of semantically consistent documents while suppressing adversarially crafted outliers. As illustrated in Figure 2, our method significantly enhances the robustness of RAG-based LLMs, mitigating adversarial influences while preserving the integrity of benign retrieval results.
We conducted comprehensive experiments on Natural Questions (NQ), MS-MARCO, and HotpotQA across five different models. Our method has shown at least a 30% decrease in reducing the Attack Success Rate (ASR), with improvements of up to 80% across various adversarial attack strategies.
We summarize our contributions as follows:
  • We introduce GRADA, a weighted similarity graph among retrieved documents iteratively propagates scores to mitigate the impacts of adversarial passages.
  • We introduce a novel scoring function that simultaneously captures both query-document and document-document correlations, thereby improving robustness against adversarial attempts to mimic the query.
  • Conducted comprehensive experiments across three different datasets against four chosen attacks. Showing GRADA’s advantages over the current defense baselines.
2 Related Work
The IR community has long examined how malicious actors distort ranking functions. Gyongyi and Garcia-Molina (2005) propose a taxonomy of web-spam strategies, categorizing attacks into content-based, link-based, and behavior-based manipulations. Content-based spam involves manipulating the textual content of web pages to boost rankings, such as stuffing keywords or inserting irrelevant but popular terms. Link-based spam, including techniques like link farms, exploits the graph structure of the web to artificially inflate a page’s authority. Ntoulas et al. (2006) focus on detecting such content-based spam using statistical features extracted from term distributions and document structure. Their work highlights how spam pages often exhibit unnatural language models and disproportionately high keyword densities. Castillo and Davison (2011) provide a comprehensive survey of adversarial techniques in web search, including cloaking (serving different content to users and crawlers), redirection, and deceptive anchor texts. These traditional adversarial IR strategies, although originally developed to mislead search engines, reveal core vulnerabilities in retrieval systems—vulnerabilities that persist, albeit in more complex forms, in modern neural IR and RAG systems.
When RAG systems came out, Corpus poisoning attacks Zhong et al. (2023) show a possible new attack surface on LLMs. However, this method does not directly affect the accuracy of the LLM; instead, it focuses on the retriever. Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al. (2023); Pasquini et al. (2024). However, compared to the prior work, these methods are unstable in ensuring the retriever retrieves the adversarial passage every time. While these attacks are situated in modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
More recently, PoisonedRAG Zou et al. (2024) was proposed as a more stable attack. It uses two passages concatenated together, with one of them appended to guarantee the retrieval of the adversarial passage and one to achieve a given adversarial goal on the generator. The goal is to let the LLM output the answer the attacker wants. PoisonedRAG inspired a lot of the new attacks. Phantom Chaudhari et al. (2024), which introduces a trigger to the question and achieves the adversarial goal only when the trigger is shown in the query. Another Prompt Injection Attack (PIA, Clop and Teglia, 2024) makes use of the passage that guarantees the retrieval in PoisonedRAG and focuses on the adversarial goal beyond misinformation.
A recent study proposed a defense mechanism that generates responses independently and produces an output based on the majority vote Xiang et al. (2024). However, this method initiates its defense at the generator stage, which can impact the accuracy of the system, especially when multiple documents are required. GRADA addresses this issue by focusing on the stage before generation, specifically the reranking process.
Figure 3: BM25 similarity matrix among retrieved documents, where D0-D4 are poisoned, and D5-D10 are clean.
3 GRADA
A defining characteristic of recent poisoning attacks on RAG Zou et al. (2024) is their focus on ensuring semantic similarity to the query while introducing anomalous similarity patterns among poisoned documents. Adversarial documents are very similar to the query but less to the legitimate documents, forming isolated patterns within the retrieval set, as illustratedFigures 2 and 3. Graph structures naturally capture these complex inter-document relationships by representing documents as nodes and similarities as edges. Leveraging this intuition, we propose a graph-based reranking method that utilizes document-document similarity to enhance retrieval robustness. In Section 3.1, we detail the graph construction process, followed by a description of our reranking system in Section 3.2.
3.1 Graph Construction
We construct a weighted, undirected graph $G=(V,E)$, where each node $v_{i}\in V$ corresponds to a document $doc_{i}$, and each edge $e_{ij}\in E$ is an undirected edge connecting node $v_{i}$ and $v_{j}$. Each edge is assigned a weight $w_{ij}\in\mathbb{R}^{+}$, which quantifies the similarity between the corresponding documents, i.e., $\text{sim}(v_{i},v_{j})$. The graph is undirected because document relationships are not inherently directional; rather, the connectivity structure defines their associations. The edge weight $w_{ij}$ can be computed using two different approaches:
  • Doc-to-Doc Similarity (D2DSIM): The weight is directly determined by the similarity between documents.
  • Hybrid Relevance Similarity (HRSIM): A function $f$ that integrates both document-document similarity and query-document relevance:
    $w_{ij}=f\big{(}\text{sim}(v_{i},v_{j}),\text{sim}(v_{i},q),\text{sim}(v_{j},q)\big{)}$
The second approach assigns edge weights that not only reflect direct document-to-document similarity but also incorporate each document’s relevance to an external query. This dual consideration leads to a more nuanced representation of document relationships.
To mitigate the influence of adversarial passages—documents that mimic the query $q$ to gain higher rankings—we introduce a function $f$, which adjusts the similarity score by applying a penalty based on the document-to-query similarities. First, we define the combined query relevance for a pair of documents $v_{i}$ and $v_{j}$ as follows:
$\text{sim}_{\text{sum}}=\text{sim}(v_{i},q)+\text{sim}(v_{j},q)$
Then, the edge weight $w_{ij}$ between $v_{i}$ and $v_{j}$ is computed by subtracting a penalty term from their direct similarity, ensuring that the weight remains non-negative:
$w_{ij}=\max\left(\text{sim}(v_{i},v_{j})-\alpha\cdot\text{sim}_{\text{sum}},0\right)$
Here, $\alpha$ is a penalty coefficient that controls the influence of query similarity. If $\text{sim}(v_{i},v_{j})<\alpha\cdot\left[\text{sim}\left(v_{i},q\right)+\text{sim}\left(v_{i},q\right)\right]$, the edge weight is set to zero, effectively removing the connection between $v_{i}$ and $v_{j}$.
Regarding the similarity function, we explore two popular methods:
  • BM25: we use BM25 Robertson and Zaragoza (2009) to calculate $\text{sim}(v_{i},v_{j})$. Since BM25 is an asymmetric metric, we adopt the following approach to compute the similarity score, ensuring symmetry in the process:
    $w_{ij}=\frac{1}{2}\left(\text{BM25}\left(v_{i},v_{j}\right)+\text{BM25}\left(v_{j},v_{i}\right)\right)$
  • Embedding-based Distance (EBD): we transform the documents $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ into dense vectors $v_{i}$ and $v_{j}$ and compute their cosine distance:
    $w_{ij}=\text{sim}(v_{i},v_{j})=\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|}$
3.2 Reranking
Inspired by PageRank Page et al. (1999), we refine document rankings through an iterative score propagation process after constructing the graph. This approach prioritizes well-connected nodes while mitigating the influence of adversarial documents, ensuring a more robust and reliable ranking.
Initially, each node $v_{i}$ is assigned a score $s^{*}_{i}$, forming the initial score vector $\mathbf{s}^{*}=[s^{*}_{1},s^{*}_{2},\ldots,s^{*}_{n}]^{\top}$. The scores are then iteratively updated at each step $t$ via:
$s_{i}^{(t)}$ $=(1-d)s^{*}_{j}+d\sum_{v_{j}\in\mathcal{N}(i)}\frac{w_{ij}}{\sum_{v_{k}\in\mathcal{N}(j)}w_{jk}}s_{j}^{(t-1)}$ (1)
where $N(i)$ represents the set of neighbor nodes connected by $v_{i}$ and $d$ is the damping factor, typically set to 0.85. The initial score vector $s^{*}$ is set by uniform initialization $s^{*}=\left[\frac{1}{|V|},\frac{1}{|V|},...,\frac{1}{|V|}\right]$.
For experiments comparing different initialization methods, please refer to Section C.2.
The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval and $M\geq n$. We retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
By ensuring that poisoned documents do not form the majority in the retrieved set (with $M\geq 2n$), we prevent adversarial documents, which may exploit the query for high relevance scores, from dominating. For example, if the original set of $n$ documents contains all poisoned ones (e.g., $n=5$), adding $\geq n$ benign documents ensures the majority is non-poisoned. This strategy guarantees that non-poisoned documents remain a significant portion of the final selection, enhancing the system’s robustness against adversarial manipulation.
After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded. These top $n$ documents are then provided as the context of the model.
4 Experiments
This section begins by detailing the experimental setup, followed by a comparison of our approach with multiple baseline methods. Finally, we analyze the effectiveness of our approach across different settings.
4.1 Experimental Setup
Attack setup.
We conduct experiments on three widely used english datasets: Natural Question Kwiatkowski et al. (2019), MS-MARCO Nguyen et al. (2016) and HotpotQA Yang et al. (2018). The victim models chosen for this study are GPT-3.5-Turbo (version 0125) Brown et al. (2020), GPT-4o (version 2024-08-06) OpenAI et al. (2024), Qwen2.5 Qwen et al. (2025) and LLaMA-3 Grattafiori et al. (2024). The prompts used to generate answers are detailed in Appendix A. Contriever Izacard et al. (2021), is a dense retriever model used to find relevant documents by calculating similarity scores between the query and the knowledge base. It was selected for this study due to its efficiency and ability to handle large datasets. In this work, we investigate four distinct attack strategies on RAG. Two of them are Black-box attacks that have no knowledge about the retriever: PoisonedRAG Zou et al. (2024) and PIA Greshake et al. (2023); Pasquini et al. (2024); Perez and Ribeiro (2022). The remaining two are white-box attacks, in which the attacker has access to the victim’s retriever: PoisonedRAG(Hotflip) Zou et al. (2024) and Phantom Chaudhari et al. (2024)
Under default settings without any defense, as in Zou et al. (2024), we retrieve the five most similar documents from the knowledge database to serve as the context for each question. We select 100 close-ended questions from each dataset, yielding 300 questions in total per attack-defense run. Additionally, this process is repeated using 3 random seeds, meaning each attack-defense pair is evaluated on 900 questions in total.
However, in contrast to Zou et al. (2024), where five poisoned texts are generated and injected into the knowledge base, To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database. The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack. As shown in Figure 3, a similarity cluster of poisoned documents appears in the top-left corner. By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries. This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
Defense setup.
We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function. [^1] We examine other similarity functions in Section 4.3 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024). We compare GRADA against two reranking models and one defense method: HLATR Zhang et al. (2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al. (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al. (2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section 3. We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation). The top five ranked documents are subsequently provided as the context for the model to answer the query. This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question. For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
Evaluation metrics.
In our experiments, we employ Attack Success Rate (ASR) and Exact Match (EM) as metrics. ASR is defined as the ratio of successful attacks to the total number of attacks conducted. An attack is considered successful if the intended poisoned answer appears as a substring within the generated response from the model. This definition accommodates attack strategies like PIA, which aim to introduce harmful links into the output of the model, allowing for some tolerance to semantically equivalent responses. A higher ASR indicates a more successful attack. This evaluation methodology follows the approach used in previous work Zou et al. (2024).
To assess the question-answering accuracy of the models, we adopt EM score. EM requires that the predicted answer of the model matches the ground truth answer exactly. This strict criterion ensures that the response of the model is precise and follows the need for exact wording specified in the query, as outlined in Appendix A.
Defense PoisonedRAG PIA
HotpotQA NQ MS-MARCO HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-3.5-Turbo
None 59.0±1.4 / 32.3±0.5 55.7±1.2 / 33.3±1.1 46.5±1.5 / 41.0±0.0 100.0±0.0 / 0.0±0.0 98.0±0.0 / 2.0±0.0 88.0±0.0 / 7.7±0.5
HLATR 62.3±0.5 / 30.3±0.5 51.5±0.5 / 35.5±0.5 36.5±1.5 / 52.0±1.0 100.0±0.0 / 0.0±0.0 92.0±0.0 / 4.0±0.0 84.0±0.0 / 9.0±0.0
BGE-reranker 56.6±0.9 / 36.3±1.2 46.5±0.5 / 43.5±0.5 34.0±0.0 / 55.0±0.0 98.0±0.0 / 2.0±0.0 43.0±0.0 / 35.7±0.5 43.0±0.0 / 43.0±0.8
Keyword Aggregation 11.0±2.0 / 62.5±2.5 2.0±0.0 / 54.0±0.0 3.0±0.0 / 60.0±2.0 0.0±0.0 / 59.0±1.0 0.0±0.0 / 48.0±0.0 0.0±0.0 / 57.5±0.5
GRADA (D2DSIM-EBD) 48.6±1.2 / 39.0±0.8 26.1±1.0 / 50.9±1.0 29.0±1.0 / 55.0±1.0 33.0±0.0 / 42.3±0.5 2.0±0.0 / 58.3±0.5 3.0±0.0 / 70.5±0.5
GRADA (D2DSIM-BM25) 45.0±0.8 / 40.0±0.5 13.5±0.7 / 55.0±1.4 16.5±0.5 / 65.5±0.5 42.0±0.0 / 33.0±0.8 12.0±0.0 / 55.3±0.5 2.0±0.0 / 69.7±0.9
GRADA (HRSIM) 10.0±0.0 / 51.0±0.8 3.0±0.6 / 58.0±1.1 8.5±0.5 / 71.5±0.5 27.0±0.0 / 41.7±1.2 2.0±0.0 / 61.7±2.1 1.0±0.0 / 74.3±0.5
Llama3.1-8b-Instruct
None 50.7±0.5 / 37.0±0.0 49.0±0.8 / 33.0±0.8 40.7±0.5 / 40.0±0.0 88.3±0.5 / 3.0±0.0 82.0±0.0 / 8.0±0.0 69.0±0.0 / 14.0±0.0
HLATR 52.3±0.5 / 35.7±0.5 39.0±0.8 / 41.3±0.5 35.7±0.5 / 43.3±0.5 91.3±0.5 / 2.7±0.5 71.7±0.5 / 15.3±0.5 50.0±0.8 / 19.7±0.5
BGE-reranker 51.7±0.5 / 36.0±0.0 42.0±0.8 / 40.7±1.2 33.7±0.5 / 42.0±0.8 79.7±0.5 / 9.7±0.9 30.0±0.0 / 40.3±0.5 19.7±0.9 / 44.7±1.2
Keyword Aggregation 6.7±1.9 / 39.0±0.8 3.0±0.0 / 39.0±0.0 6.7±0.5 / 38.3±1.2 0.0±0.0 / 35.0±0.0 0.0±0.0 / 39.0±0.0 0.0±0.0 / 36.0±0.8
GRADA (D2DSIM-EBD) 42.0±0.0 / 36.7±0.5 24.0±0.0 / 46.7±0.5 31.7±0.5 / 40.0±0.8 30.7±0.5 / 35.3±0.90 1.0±0.0 / 55.3±0.5 2.0±0.0 / 56.0±0.0
GRADA (D2DSIM-BM25) 30.0±0.0 / 39.3±0.5 8.0±0.0 / 52.3±0.5 19.3±0.5 / 49.7±0.9 39.0±0.0 / 28.7±0.5 7.7±0.5 / 48.3±0.9 0.0±0.0 / 55.0±0.0
GRADA (HRSIM) 7.0±0.0 / 44.0±0.8 2.3±0.5 / 55.7±0.5 12.0±0.0 / 52.3±0.5 23.0±0.0 / 36.7±0.5 2.0±0.0 / 55.0±0.8 0.0±0.0 / 59.3±0.5
Table 1: ASR and EM (%) for various defense methods on the black-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct. The results of other models can be found in Tables 10, 9, 11, 13 and 12. We highlight the top-2 lowest ASR results in blue cells.
Defense PoisonedRAG(Hotflip) Phantom
HotpotQA NQ MS-MARCO HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-3.5-Turbo
None 62.0±0.8 / 29.3±0.5 55.0±0.0 / 31.5±0.5 42.5±0.5 / 47.5±0.5 99.0±0.0 / 1.0±0.0 88.7±0.5 / 5.7±0.9 67.7±1.9 / 25.7±1.7
HLATR 60.7±0.5 / 30.3±0.5 49.6±0.9 / 36.0±0.8 31.3±2.1 / 55.0±2.2 97.3±0.5 / 2.7±0.5 90.7±0.5 / 7.0±0.8 64.7±9.6 / 27.3±8.2
BGE-reranker 56.6±0.5 / 34.3±1.2 43.0±0.8 / 40.7±0.5 27.3±1.2 / 59.7±0.5 94.0±0.0 / 6.0±0.0 70.7±4.7 / 17.3±0.5 57.3±9.4 / 30.7±7.4
Keyword Aggregation 12.0±0.8 / 62.3±2.1 2.0±0.0 / 52.0±4.0 4.0±0.8 / 57.0±2.6 0.0±0.0 / 50.0±0.8 0.0±0.0 / 44.0±0.0 0.0±0.0 / 57.0±1.0
GRADA (D2DSIM-EBD) 44.7±0.9 / 39.3±1.2 14.0±3.5 / 52.7±2.5 10.7±1.2 / 69.0±0.0 60.7±0.5 / 19.7±0.5 14.0±0.0 / 45.3±0.5 13.0±0.0 / 59.0±2.2
GRADA (D2DSIM-BM25) 37.0±0.8 / 44.0±0.0 9.0±0.0 / 59.3±0.5 7.3±0.9 / 70.7±0.9 27.0±0.0 / 33.0±0.8 5.7±0.5 / 50.0±0.8 0.3±0.5 / 66.0±2.2
GRADA (HRSIM) 7.3±0.5 / 52.7±0.9 4.0±0.0 / 58.3±1.2 6.3±0.9 / 72.3±1.2 23.0±0.0 / 37.3±1.2 0.0±0.0 / 48.5±0.5 0.0±0.0 / 70.0±0.5
Llama3.1-8b-Instruct
None 53.0±2.8 / 32.7±1.2 50.0±1.4 / 30.0±2.2 49.0±0.0 / 32.0±1.6 99.7±0.5 / 0.3±0.5 89.3±2.1 / 9.3±1.2 73.0±1.6 / 20.3±1.7
HLATR 53.3±2.9 / 32.7±2.1 43.7±2.1 / 37.7±2.4 36.0±1.4 / 37.7±1.7 96.7±1.2 / 3.0±0.8 92.7±1.2 / 6.0±1.4 72.3±1.2 / 18.0±1.6
BGE-reranker 50.0±3.7 / 34.3±0.5 42.3±0.5 / 36.3±1.2 27.3±1.2 / 59.7±0.5 95.3±1.2 / 3.0±0.8 72.0±1.6 / 21.7±1.7 62.0±0.8 / 26.0±1.6
Keyword Aggregation 12.0±0.8 / 62.3±2.1 2.0±0.0 / 52.0±4.0 4.0±0.8 / 57.0±2.6 0.0±0.0 / 32.0±0.0 0.0±0.0 / 36.0±0.0 0.0±0.0 / 39.7±0.9
GRADA (D2DSIM-EBD) 39.7±2.5 / 35.7±2.6 13.0±0.0 / 50.7±2.1 14.7±1.2 / 52.3±1.9 57.7±2.6 / 22.7±2.1 10.7±1.9 / 48.7±1.2 11.3±1.2 / 51.3±1.2
GRADA (D2DSIM-BM25) 32.0±0.8 / 38.0±0.0 8.7±0.9 / 52.0±0.8 13.3±0.9 / 53.0±0.8 26.7±0.5 / 37.0±2.2 4.3±0.5 / 53.7±1.2 1.0±0.0 / 56.0±0.0
GRADA (HRSIM) 8.7±1.7 / 44.7±1.2 2.0±0.8 / 53.3±2.6 6.3±0.9 / 72.3±1.2 10.3±2.1 / 41.3±1.9 0.3±0.5 / 53.7±0.9 0.0±0.0 / 60.3±1.7
Table 2: ASR and EM (%) for various defense methods on White-box attacks.
4.2 Results and Discussions
Attacking without defense.
As shown in Table 1, including a single poisoned document in the retrieval process results in a high ASR score. For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct. PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo. These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
Effectiveness of GRADA.
The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables 1 and 2. As shown in Table 1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD. With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%. However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks. On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%. Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
The defense methods demonstrate consistent effectiveness across the NQ and MS-MARCO datasets, achieving ASR reductions of over 30% in most cases. However, performance on HotpotQA is less stable, particularly for D2DSIM-EBD and D2DSIM-BM25, which achieve only around a 10% reduction in ASR against PoisonedRAG attacks. In contrast, HRSIM maintains its effectiveness, delivering ASR reductions exceeding 30%, comparable to its performance on other datasets. This discrepancy likely stems from HotpotQA’s multi-hop reasoning requirements, which pose challenges for single-document similarity metrics.
In Table 1, HLATR and BGE-reranker exhibit limited ability to filter poisoned documents, with ASR remaining largely unchanged compared to scenarios without any defense mechanisms. Although BGE-reranker occasionally outperforms HLATR, its overall performance remains inferior to GRADA in handling adversarial cases. This discrepancy underscores a critical limitation in contemporary reranking systems, which are primarily optimized for question relevance but insufficiently equipped to address adversarial attacks with high question relevance.
Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom. Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements. However, while it reduces ASR effectively, its EM scores are lower than those of GRADA. For example, on Llama3.1-8b-Instruct in Table 1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference as some critical information may be lost during keyword extraction. This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
Similar results to those presented in Table 1 can also be observed in Table 2. Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
Defense HotpotQA NQ MS-MARCO
GPT-3.5-Turbo
No-RAG 16.3±1.7 23.7±1.3 11.7±0.5
None 64.3±0.5 58.6±1.2 76.0±0.0
HLATR 70.0±0.8 62.0±0.8 77.7±0.5
BGE-reranker 68.0±1.4 64.7±1.2 78.3±0.5
Keyword Aggregation 68.3±0.5 48.0±0.0 59.0±0.0
GRADA (D2DSIM-EBD) 64.0±0.8 61.0±0.8 74.3±0.5
GRADA (D2DSIM-BM25) 57.3±0.5 64.7±0.5 75.0±1.6
GRADA (HRSIM) 50.0±0.5 62.0±0.0 75.3±0.5
Llama3.1-8b-Instruct
No-RAG 4.3±0.5 3.0±0.0 3.7±1.2
None 56.7±0.5 50.0±0.0 55.0±0.0
HLATR 56.0±0.8 51.0±0.0 56.3±0.5
BGE-reranker 58.0±0.8 54.0±0.8 59.3±0.9
Keyword Aggregation 34.0±0.0 39.0±0.0 36.0±0.8
GRADA (D2DSIM-EBD) 52.0±1.4 54.7±0.5 58.3±0.5
GRADA (D2DSIM-BM25) 47.0±0.8 52.7±0.5 54.3±0.9
GRADA (HRSIM) 43.3±0.9 54.7±0.5 57.0±0.8
Table 3: EM scores of defense methods when presented with benign inputs.
Table 3 highlights the impact of different defense mechanisms on benign inputs. On GPT-3.5-Turbo, both HLATR and BGE-reranker demonstrate strong performance, outperforming GRADA and enhancing the model’s overall accuracy. these reranking systems yield at least a 2% improvement in EM scores, suggesting their effectiveness in mitigating noise unrelated to the posed questions.
GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 4%. Notably, the use of D2DSIM-BM25 leads to an 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall. However, on HotpotQA, HRSIM resulted in an 14% reduction in EM scores when handling benign inputs. While this trade-off is significant, it corresponds to HRSIM’s remarkable ASR reduction. Striking a balance between retrieval quality and defense robustness remains a crucial challenge for future research.
Figure 4: Distribution of poisoned document positions after applying GRADA (HRSIM) in the NQ dataset. Documents positioned below rank 5 are effectively mitigated by the ranking algorithm. Other results are showed in Figures 11, 6, 7 and 8
Keyword Aggregation has a much lower performance also in EM scores on benign input compared to GRADA. For example, in MS-MARCO, it results in 40% compared to 57% on Llama3.1-8b-Instruct and 59% compared to 75.3% on GPT-3.5-Turbo. Indeed showing the cost of discarding valuable information when facing benign documents.
Using GRADA, we demonstrate that it is possible to defend against the chosen attacks without compromising the model’s overall performance on EM. While reranking methods such as HLATR and BGE-reranker show promise in reducing noise, their limited effectiveness in countering adversarial attack noise highlights a critical gap in existing defenses. Similarly, Keyword Aggregation presents a valuable strategy for mitigating attack payloads but comes with trade-offs in EM scores.
Why GRADA works.
For the attack to be effective, the attackers must ensure that the retriever selects the poisoned documents. To achieve this, they primarily focus on making these documents resemble the queries, as most retrieval models prioritize query-document similarity when selecting relevant results. Additionally, poisoned documents typically exhibit only weak similarity to other documents in the corpus. This characteristic makes them less susceptible to detection by defense mechanisms that compare retrieved documents against one another.
GRADA exploits this by constructing a document similarity graph where documents effectively "vote" for other documents they strongly resemble semantically. Benign documents, which share common relevant content, naturally have strong mutual similarities (e.g., averaging  0.82) and thus reinforce each other. In contrast, adversarial documents—designed to mislead—tend to be more isolated, receiving fewer "votes" due to their significantly lower similarities (average  0.35) with genuine documents. As a result, GRADA promotes densely interconnected benign documents while reducing the influence of sparse adversarial content. A running example is also shown in Figure 14
Figure 5: HRSIM performance with different similarity functions selection on MSMARCO dataset. The figure illustrates the proportion of test instances in which poisoned documents remain among the top five retrieved results.
4.3 Additional Studies
Ranking distribution.
We have demonstrated the effectiveness of our approach in enhancing defense performance. To gain a deeper understanding of its impact, we further analyze how our method systematically lowers the ranking of poisoned documents. As illustrated in Figure 4, the position distribution of poisoned documents within the retrieval set shifts significantly after applying GRADA with D2DSIM-BM25. Notably, over 70% of poisoned documents are relegated beyond the top five positions, substantially reducing their influence. These findings confirm that GRADA is both robust and effective in mitigating adversarial attacks.
Selections of HRSIM.
Thus far, our focus has primarily been on utilizing BM25 for HRSIM. In this section, we explore other similarity functions for HRSIM. As shown in Figure 5, we extend our analysis by incorporating SBERT Reimers and Gurevych (2019), alongside the three previously discussed methods, to better capture document-to-document similarity. Our results indicate that both EBD and SBERT exhibit strong overall performance against PIA and PoisonedRAG attacks. In contrast, BGE-Reranker struggles to effectively filter out poisoned documents, likely due to its primary training objective of computing query-to-document similarities rather than document-to-document relationships. HRSIM, when combined with BM25, effectively minimize the presence of poisoned documents, reducing them to just 14 out of 100 test instances. This outcome underscores its remarkable effectiveness in filtering malicious content.
Figure 6: Impact of the $\alpha$ value as it increases with three metrics (ASR, number of poisoned documents, and EM) on NQ dataset with GPT-3.5-Turbo.
Impact of $\alpha$ and $M$.
As shown in Figure 6, the number of poisoned documents in the context decreases as $\alpha$ increases, reaching a minimum at $\alpha=0.3$ before starting to rise again after $\alpha=0.8$. The ASR follows a similar trend to the number of poisoned documents after $\alpha=0.3$. Conversely, the EM score exhibits a minimum at $\alpha=0.7$. We selected $\alpha=0.4$ because it strikes a balance, avoiding excessive penalization for query similarity, which could otherwise result in fewer query-related documents. When $\alpha=0.4$, all three metrics (ASR, number of poisoned documents, and EM) are within an acceptable range, approaching the optimal performance values for $\alpha$.
Figure 7: Impact of the $M$ value as it changes with three metrics (ASR, number of poisoned documents, and EM) on MSMARCO dataset with GPT-3.5-Turbo.
Defense Total Time (s) Processing Time (s) Defense-Only Time (s) Defense-Only Processing (s)
Keyword Aggregation 12.61 11.11 11.59 9.21
GRADA (D2DSIM-EBD) 1.56 1.12 0.62 0.62
GRADA (D2DSIM-BM25) 0.97 0.52 0.02 0.02
GRADA (HRSIM) 1.05 0.61 0.05 0.05
Table 4: Runtime Comparison (on GPT-3.5-Turbo, average per query): Total Time (s) and Processing Time (s) represent the complete runtime for answering one question, including retrieval, defense method, and LLM response generation. In contrast, Defense-Only Time (s) measures exclusively the runtime of the defense methods themselves. Total Time is recorded using Python’s time.time() function, whereas Processing Time is measured with Python’s time.process_time() function.
Figure 7 illustrates the effect of selecting $M=n$. It shows that, regardless of how documents are re-ranked, poisoned documents can still remain within the context provided to the model. However, this approach results in a 17% decrease in ASR and a 9% increase in EM, indicating that simply adjusting document positions can significantly impact model performance. This aligns with our observations in Table 5, and the specific positions of the documents are detailed in Figure 4. By including additional documents for reranking and then retrieving only the top $n$ results, the ASR is further reduced from 21% to 10%, with only 14% of poisoned documents remaining in the context provided to the model. This demonstrates the importance of including extra documents during reranking to remove poisoned content and achieve better overall performance effectively.
Computational Complexities.
The overall complexity of GRADA consists of two main components:
  • Similarity matrix construction: $O(N^{2})$, where N is the number of retrieved documents. This step can incur additional costs depending on the chosen similarity function. For example, using D2DSIM-EBD (embedding-based document similarity), the complexity becomes $O(N^{2}\cdot d)$, where $d$ is the embedding dimension. BM25-based similarity: the complexity is $O(N^{2}\cdot L)$, where L is the average document length. This is efficient due to the sparsity of token overlaps and inverted index optimizations. Here, since we are reranking the documents after the retrieval step. The retrieved documents set is usually constrained with limited amounts of data, making this a viable solution.
  • Graph-based reranking (e.g., PageRank): $O(n+m)$, where n is the number of nodes (documents) and m is the number of edges in the constructed similarity graph.
The only defense Keyword Aggregation requires querying the language model N times—once per document—to collect individual answers before aggregating: $O(N*C_{LM})$ (where $C_{LM}$ refers to the language model’s cost). This incurs significantly higher costs in terms of API calls and model generation time, especially with large models.
GRADA, by comparison, does not require any model calls. The only required model call is after GRADA to query the final answer, making it more efficient and scalable for large-scale or production RAG deployments.
5 Conclusion
The study examines the robustness challenges faced by RAG systems. We identify a critical vulnerability in current adversarial attacks, which focus on increasing semantic similarity to the query without accounting for the relationships between the retrieved documents. Our proposed graph-based filtering framework, GRADA, enhances the robustness of RAG systems by leveraging document similarities and effectively mitigating adversarial impacts through information flow. Experimental results on datasets such as MS-MARCO and NQ, demonstrate at least 30% reductions in ASR across various adversarial strategies. Overall, this work presents a promising direction for developing more secure and reliable RAG systems.
Limitations
Despite its effectiveness, our approach has limitations. First, it struggles with multi-hop reasoning tasks, facing attacks like PIA and Phantom. As the number of poisoned documents increases, system robustness deteriorates. Second, our method assumes poisoned documents are a minority. When they form the majority, their effectiveness declines, and future work should explore adaptive retrieval strategies to counter adversarial dominance.
Ethics Statement
Our study focuses on improving the robustness of RAG systems, thereby enhancing their reliability and minimizing harmful manipulations. We evaluated our proposed method, GRADA, using publicly available datasets as detailed in Appendix F. We do not engage in harmful data practices.
Acknowledgement
Aryo Pradipta Gema was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. Giwon Hong was supported by the ILCC PhD program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics. Xuanli He was funded by an industry grant from Cisco. Pasquale Minervini was partially funded by ELIAI, an industry grant from Cisco, and a donation from Accenture LLP. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
Appendix A Prompts to LLM
Figure 8 shows the two prompts we use to query the LLMs.
Prompts given to LLMs Prompt 1: You are a helpful assistant, below is a query from a user and some relevant contexts. Answer the question given the information in those contexts / keywords. Your answer should be short ,concise and without explanations ,just answer with one or two words like yes or no. If you cannot find the answer to the question, just say "I don’t know". Contexts / Keywords: [context / keywords] Query: [question] Answer: Prompt 2: You are a helpful assistant, below is a query from a user and some relevant contexts. Answer the question given the information in those contexts. If you cannot find the answer to the question, just say "I don’t know". Contexts: [context] Query: [question] Answer: Prompt 3: You are a helpful assistant, Answer the query with no more than ten words. If you do not know the answer confidently, just say "I don’t know". Query: [question] Answer:
Figure 8: Example of prompts given to LLM, Prompt 1 is used in all of the defense methods to generate the final output. Prompt 2 is only used in the phase to generate keywords and Prompt 3 is used for No-RAG from the paper Xiang et al. (2024).
Appendix B Deconstructing PoisonedRAG
PoisonedRAG Example Question: "When did the apple iphone se come out?" Benign documents 1: "March 21, 2017 iPhone SE shipped with iOS 9.3, supporting iPhone 6 and 6S-exclusive features such as Apple Pay, Live Photos, Retina Flash and always-on Siri voice activation.[6][10][16]" Benign documents 2: "On launch, it was released in models with either 16 or 64 GB of internal storage; on March 21, 2017, Apple announced that these models would be replaced by new improved models at the same launch price points, releasing on March 24, 2017.[11][12]" Poisoned Document 1: "When did the apple iphone se come out. The Apple archive, housing important launch events, mentions a significant one on April 15, 2015. On this day, the compact, budget-friendly iPhone SE hit the market, impressing many gadget enthusiasts." Poisoned Document 2: "When did the apple iphone se come out. The iPhone series has released several models. Notable amongst them is the iPhone SE, introduced to the market on April 15, 2015, attracting considerable consumer attention."
Figure 9: Example of PoisonedRAG attacks. Poisoned documents injected into the database are all very similar to each other and focus solely on ensuring similarity to the query, the similarities among the retrieved documents are never considered.
PoisonedRAG Zou et al. (2024) is an adversarial attack on RAG systems that operates in two stages. The first part enhances the semantic similarity of the adversarial passage to the query, increasing the likelihood of it being retrieved. The second part introduces adversarial content to mislead the model into generating a specific incorrect response.
While the approach used to achieve the first part of the attack is effective, it is also simple. Specifically, the adversarial passage is constructed by prepending the query into the poisonous passage. Despite its simplicity, PoisonedRAG degrades the accuracy of the LLMs significantly. As shown in Table 5 (first row), the attack achieves an ASR of 54.3% on average across three datasets with just one adversarial passage retrieved as the most similar to the query.
Attack Method HotpotQA NQ MS-MARCO Average
Normal retrieved 59.0 56.0 48.0 54.3
   w/o question 66.0 61.0 51.0 59.3
Poisoned in the middle 59.0 54.0 37.0 50.0
   w/o question 63.0 51.0 34.0 49.3
Table 5: PoisonedRAG Attack Success Rate (%) where the retrieval part is removed, and the poisoned documents are placed in the middle.
Our analysis reveals that the prepended query in the adversarial passage does not significantly affect the ASR. As shown in Table 5 (second row), removing the prepended query leads to an increase in the ASR. This shows that the query was prepended only to ensure that the retriever retrieves the adversarial document, but not affecting the accuracy significantly. Furthermore, Table 5 (third and fourth row) shows that the position of the poisoned document within the retrieved documents set influences the ASR significantly, with a decrease in average ASR of 10%. This phenomenon is similar to the lost-in-the-middle effect Liu et al. (2024), where the position of the document impacts its effectiveness in influencing the output of the reader model.
Due to its straightforward approach of prepending the query to the adversarial documents, PoisonedRAG attacks can be easily identified. As demonstrated in Figure 3 and Figure 9, the attacks injected into the database often exhibit considerable similarity to one another. By focusing on the similarities between the documents in the retrieved set, we can filter out adversarial passages and decrease the ASR.
(a) ASR of GRADA as poisoned documents increase.
(b) Total poisoned documents remain after filtering.
Figure 10: Impact of increasing poisoned documents on GRADA’s performance in NQ dataset (GPT-3.5-Turbo, $M=10$).
(a) D2DSIM-EBD
(b) D2DSIM-BM25
(c) HRSIM
Figure 11: Distribution of Ground Truth document positions after applying GRADA in the NQ dataset with different ranking methods.
Appendix C Ablation Study
C.1 Number of poisoned documents increase
As shown in Figure 10(a), the effectiveness of GRADA reduces as the proportion of poisoned documents increases. When using D2DSIM-EBD, the ASR achieved by GRADA approaches that of an undefended system. However, HRSIM remains effective, achieving a 27% reduction in ASR even when half of the retrieved documents are adversarial. This is further supported by Figure 10(b), which shows that 38% of poisoned documents are still successfully filtered.
C.2 Different initial score vector
Different initial score vectors can have a significant impact on the final distribution of documents in certain cases. For instance, we experimented with initializing the score vector with query-document similarity $s^{*}=\left[\frac{sim(q,v_{0})}{\sum_{j=0}^{n}sim(q,v_{j})},\frac{sim(q,v_{1})}{\sum_{j=0}^{n}sim(q,v_{j})},...,\frac{sim(q,v_{n})}{\sum_{j=0}^{n}sim(q,v_{j})}\right]$. As shown in Figure 12(a), using a query-document initialization results in more documents being positioned between rank 5 and 8, rather than lower. We hypothesize that this is because adversarial documents may receive disproportionately high initial scores compared to benign documents. Such an imbalance gives adversarial documents a substantial advantage, particularly when the edge weights between documents are relatively small. In these scenarios, the graph-based reranking process may struggle to compensate for this initial disparity, as illustrated in Figure 14. From the analysis in Figure 12(b), we observe that this phenomenon is more prevalent in datasets like HotpotQA.
(a) Distribution of Poisoned document positions after applying GRADA (HRSIM) with different initialization in the NQ dataset.
(b) Total number of poisoned documents after applying GRADA (HRSIM) with different initialization in the NQ dataset.
Figure 12: Impact of different initialization score vectors on GRADA’s performance ($M=10$).
Appendix D Different initial score vector demonstration
Initial Score Example in Figure 14 Question: "Are Random House Tower and 888 7th Avenue both used for real estate?" Documents 1: "The former Bertelsmann Building, now known as 1540 Broadway, is a 44-story, 733 foot (223 m) office tower in Times Square in Manhattan…" Documents 2: "The Random House Tower, also known as the Park Imperial Apartments, is a 52-story mixed-use tower in New York City, United States, that is…" Documents 3: "888 7th Avenue is a 628 ft (191m) tall modern-style office skyscraper in Midtown Manhattan which was completed in 1969 and has 46 floors. Emery Roth & Sons designed…" Documents 4: "What do the estates of film stars Vincent Price and Glenn Ford have in common? And what do each of these estates have in common with valuables owned by Laugh-In’s Arte…" Documents 5: "750 Seventh Avenue is a 615 ft (187m) tall Class-A office skyscraper in New York City. It was completed in 1989 in the postmodern style and has 36 floors…" Documents 6: "The Fisk Towers is a front for the Kingpin (Wilson Fisk)’s public ventures as well as a base of operations for his criminal activities, until…" Document 0: "Are Random House Tower and 888 7th Avenue both used for real estate?.Random House Tower is occupied by a publishing company, not devoted to real estate. 888 7th Avenue is primarily used for law firms, again not real estate operations."
Figure 13: Document examples used to generate Figure 14 to demonstrate different initial score vector and their results when the adversarial documents receive significantly higher initial scores compared to benign documents. Red Documents indicates the poisoned document.
Figure 13 shows the documents we used in Figure 14.
Figure 14: A demonstration on different initial score vector and their results when the adversarial documents receive significantly higher initial scores compared to benign documents. This is an example from the HotpotQA dataset with the question:" Are Random House Tower and 888 7th Avenue both used for real estate?". The top 4 ranked documents are listed with bold final values.
Appendix E Computational Resources
The cost of a single defense run on GPT-3.5-Turbo is \$0.50, identical to a standard query since the method does not introduce additional API calls. Experiments for LLaMA-3 and Qwen2.5 were conducted on a single NVIDIA A100 80GB GPU, with each defense run taking one hour to complete.
Appendix F License and Distribution Terms
The dataset used in our experiments is publicly available under Creative Commons Attribution 4.0 International (MS-MARCO) and Apache License 2.0 (NQ, HotpotQA). The code used in our experiments is publicly available under MIT License (BM25s, PoisonedRAG) https://github.com/agrzheng/GRADA
Defense Method PoisonedRAG PoisonedRAG(Hotflip) PIA Phantom
No Defense 99.0 99.0 96.0 94.0
HLATR 100.0 100.0 93.0 89.0
BGE-reranker 100.0 98.0 47.0 58.0
GRADA (D2DSIM-EBD) 55.0 20.0 6.0 5.0
GRADA (D2DSIM-BM25) 25.0 16.0 6.0 4.0
GRADA (HRSIM) 13.0 8.0 7.0 2.0
Table 6: The percentage of poisoned documents in the given context to LLM before and after different defense methods on the NQ dataset. A lower value is better. Method Keyword not included as it does not conduct reranking.
Defense Method PoisonedRAG PoisonedRAG(Hotflip) PIA Phantom
No Defense 98.0 99.0 89.0 65.0
HLATR 98.0 96.0 85.0 70.0
BGE-reranker 98.0 98.0 48.0 53.0
GRADA (D2DSIM-EBD) 69.0 22.0 10.0 10.0
GRADA (D2DSIM-BM25) 34.0 15.0 2.0 2.0
GRADA (HRSIM) 19.0 8.0 1.0 2.0
Table 7: The percentage of poisoned documents in the given context to LLM before and after different defense methods on the MS-MARCO dataset. A lower value is better. Method Keyword not included as it does not conduct reranking.
Defense Method PoisonedRAG PoisonedRAG(Hotflip) PIA Phantom
No Defense 100.0 100.0 100.0 100.0
HLATR 100.0 100.0 100.0 99.0
BGE-reranker 98.0 100.0 98.0 98.0
GRADA (D2DSIM-EBD) 84.0 66.0 52.0 49.0
GRADA (D2DSIM-BM25) 64.0 53.0 35.0 32.0
GRADA (HRSIM) 19.0 18.0 26.0 20.0
Table 8: The percentage of poisoned documents in the given context to LLM before and after different defense methods on the HotpotQA dataset. A lower value is better. Method Keyword not included as it does not conduct reranking.
Model Defense HotpotQA NQ MS-MARCO
GPT-4o No-RAG 26.0±0.0 20.0±0.0 14.0±0.0
None 60.7±1.2 58.7±0.5 65.0±0.0
HLATR 63.3±2.6 61.3±0.5 67.3±0.5
BGE-reranker 62.7±1.2 66.0±0.8 70.3±1.9
Keyword Aggregation 61.7±0.9 47.0±0.0 47.0±0.0
GRADA (D2DSIM-EBD) 57.3±0.9 55.0±0.8 62.0±0.0
GRADA (D2DSIM-BM25) 56.3±0.9 58.7±0.5 66.0±0.0
GRADA (HRSIM) 51.0±1.4 62.0±0.0 65.0±0.0
Llama3.1-70b-Instruct No-RAG 26.0±1.6 21.3±1.2 15.3±0.5
None 60.0±0.8 66.7±0.5 53.7±0.5
HLATR 62.7±0.5 60.7±0.5 53.0±0.8
BGE-reranker 59.0±0.0 66.0±0.0 55.7±0.5
Keyword Aggregation 24.3±1.20 22.3±0.5 15.3±0.5
GRADA (D2DSIM-EBD) 47.7±0.5 55.7±0.5 46.3±0.5
GRADA (D2DSIM-BM25) 39.0±0.8 58.3±0.5 52.7±0.5
GRADA (HRSIM) 32.0±0.8 54.7±0.9 53.0±0.8
Qwen2.5-7b-Instruct No-RAG 6.0±0.0 11.0±0.0 4.7±0.5
None 46.0±0.0 50.7±1.2 51.0±0.0
HLATR 47.3±0.5 48.0±0.0 44.0±1.4
BGE-reranker 44.7±0.5 49.0±0.0 47.3±0.5
Keyword Aggregation 13.0±0.0 16.0±0.0 23.0±0.8
GRADA (D2DSIM-EBD) 40.7±0.5 45.7±0.5 44.0±0.8
GRADA (D2DSIM-BM25) 37.7±0.5 48.3±0.5 46.0±0.8
GRADA (HRSIM) 30.0±0.0 44.7±1.2 48.7±0.9
Table 9: EM scores of defense methods on GPT-4o, Llama3.1-70b-Instruct and Qwen2.5-7b-Instruct when presented with benign inputs.
Model Defense HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-4o None 42.0±0.0 / 40.0±0.8 29.3±0.9 / 38.0±1.4 24.0±0.0 / 46.0±0.0
HLATR 38.0±0.8 / 44.7±2.0 27.0±0.8 / 48.3±1.2 21.0±0.0 / 53.0±0.0
BGE-reranker 37.3±1.2 / 45.7±1.7 24.7±0.5 / 54.7±0.5 20.0±0.0 / 54.0±0.0
Keyword Aggregation 6.7±0.5 / 58.3±1.9 1.0±0.0 / 46.0±0.0 5.7±0.5 / 45.7±0.5
GRADA (D2DSIM-EBD) 37.0±0.0 / 42.0±0.8 9.7±0.5 / 46.3±0.9 19.0±0.0 / 51.0±0.0
GRADA (D2DSIM-BM25) 23.7±0.5 / 43.3±0.9 5.0±0.0 / 60.0±0.0 10.0±0.0 / 64.0±0.0
GRADA (HRSIM) 5.0±0.0 / 50.0±1.4 1.0±0.0 / 64.0±1.4 4.0±0.0 / 67.0±0.0
Llama3.1-70b-Instruct None 57.7±0.9 / 37.3±0.9 56.7±0.5 / 29.7±0.5 54.3±1.2 / 28.3±0.9
HLATR 53.0±0.8 / 43.3±0.5 49.0±0.0 / 39.0±0.0 40.7±1.2 / 37.0±0.8
BGE-reranker 53.3±0.5 / 41.7±0.5 49.3±0.5 / 38.0±0.0 37.0±0.8 / 37.0±0.8
Keyword Aggregation 4.7±0.5 / 26.0±0.8 3.0±0.0 / 22.3±0.5 3.0±0.0 / 58.0±2.2
GRADA (D2DSIM-EBD) 45.3±0.5 / 37.3±0.5 26.0±0.0 / 44.0±0.0 34.3±0.5 / 39.3±0.5
GRADA (D2DSIM-BM25) 36.0±0.0 / 37.7±0.5 11.0±0.0 / 56.3±0.5 15.0±0.0 / 50.7±0.5
GRADA (HRSIM) 8.3±0.5 / 37.7±0.5 2.7±0.5 / 53.0±0.0 9.0±0.0 / 54.0±0.0
Qwen2.5-7b-Instruct None 62.0±0.0 / 24.0±0.0 50.3±0.5 / 26.7±0.5 49.0±0.0 / 28.7±0.5
HLATR 60.0±0.0 / 28.7±0.5 42.7±0.5 / 31.3±0.5 41.0±0.0 / 28.0±0.8
BGE-reranker 60.0±0.0 / 30.7±0.5 47.7±0.5 / 29.3±0.5 42.0±0.8 / 29.3±0.5
Keyword Aggregation 4.7±0.5 / 6.0±0.0 3.0±0.0 / 11.0±0.0 9.3±0.9 / 25.0±0.8
GRADA (D2DSIM-EBD) 57.0±0.0 / 24.3±0.5 24.3±0.5 / 35.3±1.2 37.3±0.5 / 31.3±0.5
GRADA (D2DSIM-BM25) 42.3±0.5 / 27.7±0.5 12.7±0.5 / 45.0±0.8 23.7±0.5 / 38.0±1.4
GRADA (HRSIM) 7.7±0.5 / 34.0±0.8 5.3±0.5 / 41.0±0.0 12.3±0.5 / 39.0±0.8
Table 10: ASR and EM (%) for various defense methods on PoisonedRAG on GPT-4o, Llama3.1-70b-Instruct and Qwen2.5-7b-Instruct. Blue cells indicate top-two lowest ASR.
Model Defense HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-4o None 45.3±0.5 / 41.7±0.5 32.3±0.5 / 39.0±1.4 24.7±0.9 / 46.0±1.4
HLATR 42.0±0.8 / 45.0±0.8 28.3±0.9 / 48.7±1.9 19.7±2.4 / 53.0±1.4
BGE-reranker 40.0±0.0 / 41.3±0.5 27.0±0.0 / 49.0±0.0 20.0±0.0 / 53.7±0.9
Keyword Aggregation 8.7±0.5 / 59.3±1.9 1.0±0.0 / 46.0±0.0 4.0±0.0 / 48.0±1.4
GRADA (D2DSIM-EBD) 31.7±0.5 / 45.3±1.2 5.0±0.8 / 55.3±1.2 11.3±0.5 / 56.0±1.4
GRADA (D2DSIM-BM25) 21.0±0.0 / 46.3±0.9 5.0±0.0 / 61.3±0.5 7.3±0.5 / 67.0±1.4
GRADA (HRSIM) 5.0±0.0 / 49.3±1.9 1.0±0.0 / 63.3±2.4 4.0±0.0 / 66.3±0.5
Llama3.1-70b-Instruct None 56.7±0.9 / 33.3±0.9 54.7±2.1 / 26.7±1.7 47.7±0.5 / 29.0±0.8
HLATR 52.0±2.2 / 37.3±1.2 47.3±2.1 / 35.7±0.9 32.3±0.5 / 37.0±0.8
BGE-reranker 48.3±1.2 / 44.3±1.9 42.7±1.2 / 41.3±1.2 35.7±1.9 / 33.3±0.9
Keyword  Xiang et al. (2024) 4.7±0.5 / 26.0±0.8 3.0±0.0 / 22.0±0.0 3.0±0.0 / 57.0±0.8
GRADA (D2DSIM-EBD) 37.0±0.0 / 40.3±1.7 11.0±1.6 / 48.7±2.1 15.3±1.7 / 45.7±0.5
GRADA (D2DSIM-BM25) 33.3±0.9 / 37.7±2.1 6.7±0.5 / 56.0±0.8 10.3±0.5 / 51.7±1.2
GRADA (HRSIM) 8.7±0.5 / 36.7±2.6 1.0±0.0 / 54.0±0.8 6.7±0.5 / 52.3±2.4
Qwen2.5-7b-Instruct None 58.7±0.9 / 30.7±1.2 58.0±2.2 / 22.3±2.1 51.0±1.4 / 31.3±2.9
HLATR 55.7±0.9 / 33.7±2.1 51.0±0.0 / 29.0±0.8 36.3±3.3 / 33.0±3.3
BGE-reranker 54.0±1.6 / 33.7±2.6 51.0±0.8 / 29.3±0.5 37.3±4.0 / 33.3±3.9
Keyword 4.7±0.5 / 6.0±0.0 3.0±0.0 / 11.0±0.0 10.3±0.5 / 23.7±0.5
GRADA (D2DSIM-EBD) 45.7±0.9 / 31.0±1.6 14.7±1.7 / 41.0±3.6 19.0±1.6 / 36.3±0.5
GRADA (D2DSIM-BM25) 38.3±0.5 / 31.7±1.2 12.0±2.2 / 42.0±0.8 14.7±1.2 / 40.7±0.5
GRADA (HRSIM) 6.0±0.0 / 33.0±0.0 4.3±0.9 / 45.3±0.5 10.7±0.5 / 39.0±1.4
Table 11: ASR and EM (%) for various defense methods on PoisonedRAG(Hotflip). Blue cells indicate top-two lowest ASR.
Model Defense HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-4o None 99.0±0.0 / 0.3±0.5 95.7±0.5 / 3.7±0.5 80.0±0.0 / 11.0±0.0
HLATR 97.6±0.9 / 1.3±0.9 78.0±0.0 / 15.0±0.0 53.0±0.0 / 32.0±0.0
BGE-reranker 87.3±0.5 / 7.0±1.4 36.0±1.4 / 39.7±0.9 24.0±0.0 / 51.0±0.0
Keyword Aggregation 0.0±0.0 / 53.7±2.4 0.0±0.0 / 44.0±0.0 0.0±0.0 / 45.7±0.5
GRADA (D2DSIM-EBD) 30.7±0.5 / 42.3±0.9 2.0±0.0 / 57.3±0.5 2.0±0.0 / 60.0±0.0
GRADA (D2DSIM-BM25) 40.0±1.4 / 36.3±0.9 10.7±0.9 / 57.3±0.9 0.0±0.0 / 68.0±0.0
GRADA (HRSIM) 25.0±0.0 / 42.7±0.5 1.0±0.0 / 63.7±0.9 0.0±0.0 / 68.0±0.0
Llama3.1-70b-Instruct None 100.0±0.0 / 0.0±0.0 98.0±0.0 / 2.0±0.0 88.0±0.0 / 8.0±0.0
HLATR 100.0±0.0 / 0.0±0.0 91.7±0.5 / 5.3±0.5 84.0±0.0 / 8.7±0.5
BGE-reranker 98.0±0.0 / 2.0±0.0 42.3±0.5 / 38.7±0.5 43.0±0.0 / 30.3±0.5
Keyword Aggregation 0.0±0.0 / 26.7±0.5 0.0±0.0 / 23.0±1.4 0.0±0.0 / 59.3±0.9
GRADA (D2DSIM-EBD) 33.0±0.0 / 29.0±0.0 2.0±0.0 / 55.3±0.5 3.0±0.0 / 49.0±0.8
GRADA (D2DSIM-BM25) 42.0±0.0 / 25.0±0.0 12.0±0.0 / 52.0±0.8 2.0±0.0 / 54.3±1.2
GRADA (HRSIM) 26.0±0.0 / 32.0±0.8 1.3±0.5 / 55.3±0.5 1.0±0.0 / 54.7±0.5
Qwen2.5-7b-Instruct None 5.3±0.5 / 22.7±0.5 5.7±0.5 / 17.0±0.0 6.0±0.0 / 27.0±0.8
HLATR 14.0±0.8 / 24.0±1.4 17.7±0.9 / 12.7±0.9 18.0±0.0 / 20.7±0.5
BGE-reranker 25.0±0.0 / 17.0±0.0 23.0±0.0 / 31.7±0.5 18.3±0.5 / 32.0±0.0
Keyword Aggregation 0.0±0.0 / 6.0±0.0 0.0±0.0 / 11.0±0.0 0.0±0.0 / 21.3±0.5
GRADA (D2DSIM-EBD) 12.0±0.0 / 34.7±0.9 2.0±0.0 / 47.0±0.8 3.0±0.0 / 41.7±0.5
GRADA (D2DSIM-BM25) 15.0±0.0 / 28.0±0.0 8.0±0.0 / 43.7±0.5 1.0±0.0 / 44.0±1.4
GRADA (HRSIM) 8.7±0.5 / 35.3±0.5 2.0±0.0 / 46.3±0.9 1.0±0.0 / 47.3±1.7
Table 12: ASR and EM (%) for various defense methods on PIA. Blue cells indicate top-two lowest ASR.
Model Defense HotpotQA NQ MS-MARCO
ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
GPT-4o None 57.3±0.5 / 25.3±0.9 37.0±0.0 / 18.7±0.5 21.0±0.0 / 45.0±0.0
HLATR 47.0±1.4 / 27.3±0.5 36.3±0.5 / 22.3±0.5 18.0±0.0 / 53.0±0.0
BGE-reranker 35.7±0.9 / 29.7±0.5 20.3±0.5 / 32.3±0.5 19.0±0.0 / 53.0±0.0
Keyword Aggregation 0.0±0.0 / 57.0±0.0 0.0±0.0 / 48.0±0.0 0.0±0.0 / 45.0±0.0
GRADA (D2DSIM-EBD) 30.0±1.4 / 35.3±0.5 3.7±0.5 / 43.3±1.9 2.0±0.0 / 53.0±0.0
GRADA (D2DSIM-BM25) 7.3±0.9 / 40.0±1.4 2.0±0.0 / 51.0±0.0 0.3±0.5 / 63.0±0.0
GRADA (HRSIM) 3.3±0.9 / 41.3±0.9 0.0±0.0 / 50.0±1.4 0.0±0.0 / 63.7±0.5
Llama3.1-70b-Instruct None 98.7±0.5 / 1.3±0.5 90.7±1.2 / 7.3±1.2 74.3±1.2 / 19.7±0.5
HLATR 98.0±0.8 / 0.7±0.5 93.7±0.9 / 5.3±0.5 78.0±1.6 / 13.3±0.9
BGE-reranker 96.3±0.5 / 3.7±0.5 75.7±0.9 / 14.0±0.8 70.7±0.9 / 20.3±1.7
Keyword Aggregation 0.0±0.0 / 18.7±0.5 0.0±0.0 / 17.3±0.5 0.0±0.0 / 51.3±1.2
GRADA (D2DSIM-EBD) 60.3±2.9 / 16.3±1.2 12.7±2.6 / 41.7±1.7 13.7±2.4 / 45.3±2.1
GRADA (D2DSIM-BM25) 27.0±1.4 / 27.7±1.2 5.3±0.5 / 49.3±0.5 1.3±0.5 / 55.3±0.5
GRADA (HRSIM) 11.3±0.9 / 27.3±1.2 0.7±0.5 / 50.7±1.2 0.0±0.0 / 56.0±0.8
Qwen2.5-7b-Instruct None 58.7±3.8 / 18.3±1.2 56.0±2.9 / 12.0±2.2 40.0±1.4 / 25.3±2.1
HLATR 63.0±1.4 / 17.7±2.1 71.0±1.6 / 9.3±1.7 48.3±2.1 / 18.7±2.1
BGE-reranker 62.3±4.1 / 19.7±0.5 57.7±2.6 / 19.7±1.2 50.3±0.9 / 25.7±2.5
Keyword Aggregation 0.0±0.0 / 1.0±0.0 0.0±0.0 / 5.0±0.0 0.0±0.0 / 5.0±0.0
GRADA (D2DSIM-EBD) 41.0±2.8 / 17.0±3.7 11.0±2.8 / 32.0±0.8 11.7±1.7 / 40.7±2.1
GRADA (D2DSIM-BM25) 24.0±0.0 / 27.7±2.1 5.3±0.5 / 35.3±1.2 0.3±0.5 / 45.7±0.9
GRADA (HRSIM) 14.0±2.4 / 27.3±0.9 0.7±0.5 / 36.3±0.5 0.0±0.0 / 48.7±1.7
Table 13: ASR and EM (%) for various defense methods on Phantom. Blue cells indicate top-two lowest ASR.

### v2

@@
-GRADA: Graph-based Reranker against Adversarial Documents Attack
+GRADA: Graph-based Reranking against Adversarial Documents Attack
 Abstract
-Retrieval Augmented Generation (RAG) frameworks improve the accuracy of large language models (LLMs) by integrating external knowledge from retrieved documents, thereby overcoming the limitations of models’ static intrinsic knowledge.
+Retrieval Augmented Generation (RAG) frameworks can improve the factual accuracy of large language models (LLMs) by integrating external knowledge from retrieved documents, which is useful for overcoming the limitations of models’ static intrinsic knowledge.
 However, these systems are susceptible to adversarial attacks that manipulate the retrieval process by introducing documents that are adversarial yet semantically similar to the query.
 Notably, while these adversarial documents resemble the query, they exhibit weak similarity to benign documents in the retrieval set.
-Thus, we propose a simple yet effective Graph-based Reranking against Adversarial Document Attacks (GRADA) framework aiming at preserving retrieval quality while significantly reducing the success of adversaries.
-Our study evaluates the effectiveness of our approach through experiments conducted on five LLMs: GPT-3.5-Turbo, GPT-4o, Llama3.1-8b-Instruct, Llama3.1-70b-Instruct, and Qwen2.5-7b-Instruct.
-We use three datasets to assess performance, with results from the Natural Questions dataset demonstrating up to an 80% reduction in attack success rates while maintaining minimal loss in accuracy.
-GRADA: Graph-based Reranker against Adversarial Documents Attack
-Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3, Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu5 1University of Melbourne, 2University of Edinburgh, 3University College London 4University of Manchester, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@manchester.ac.uk, qiongkai.xu@mq.edu.au
+Thus, we propose a simple yet effective Graph-based Reranking against Adversarial Document Attacks (GRADA) framework aimed at preserving retrieval quality while significantly reducing the success of adversaries.
+Our study evaluates the effectiveness of our approach through experiments conducted on six LLMs: GPT-3.5-Turbo, GPT-4o, Llama3.1-8b-Instruct, Llama3.1-70b-Instruct, Qwen2.5-7b-Instruct, and Qwen2.5-14b-Instruct.
+We use three datasets to assess performance, with results from the Natural Questions dataset showing up to an 80% reduction in attack success rates while maintaining minimal loss in accuracy.
+GRADA: Graph-based Reranking against Adversarial Documents Attack
+Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Author., Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
 1 Introduction
@@
 (2024), subtly altering rankings to mislead LLM outputs.
-As shown in Figure 1, these adversarial documents mimic query-relevant patterns, making them difficult to detect while degrading the reliability of retrieval-based LLM systems.
-In practice, LLMs are used in the search engine to produce direct answers to user’s query, called Answer Engine Optimization (AEO)  Yalçın and Köse (2024).
-Adversarial documents can be created using the previous methods to achieve higher rankings in the search results.
-Also, guide the search engine to produce malicious context that can potentially harm the user Hammond (2024); Venkit et al.
+As shown in Figure˜1, these adversarial documents mimic query-relevant patterns, making them difficult to detect while degrading the reliability of retrieval-based LLM systems.
+In real-world applications, LLMs are increasingly used in search engines to provide direct answers to user queries, a process known as Answer Engine Optimization (AEO) Yalçın and Köse (2024).
+By leveraging retrieval-time manipulation techniques, attackers can craft adversarial content that not only ranks higher in search results but also steers the generated answers toward harmful or misleading content Hammond (2024); Venkit et al.
 (2024).
 Figure 2: An overview of GRADA.
-A vanilla RAG pipeline concatenates all retrieved documents along with the question as the input to the LLM.
+A vanilla RAG pipeline concatenates all retrieved documents along with the question as input to the LLM.
 However, the accuracy of this pipeline can be easily harmed by malicious passages.
-In contrast, GRADA uses a graph-based approach to isolate and filter out malicious passages before passing the retrieved documents as the LLM input.
+In contrast, GRADA uses a graph-based approach to rerank and filter out malicious passages before passing the retrieved documents to LLMs for generation.
 Existing noise filtering methods, such as Hybrid List Aware Transformer Reranking (HLATR, Zhang et al., 2022) and BAAI General Embeddings (BGE-reranker, Xiao et al., 2023), focus on improving document relevance by filtering out generic noise or low-quality content.
 However, these methods are ineffective against adversarial attacks that exploit query-document similarity patterns to evade detection.
-In addition, a recent study has reviewed current graph-based reranking methods.
-It shows a potential path of using graphs in future information retrieval Zaoad et al.
-(2025) tasks but the effects on adversarial documents remains unknown.
+In addition, a recent study has reviewed current graph-based reranking methods Zaoad et al.
+(2025).
+It shows a potential path to use graphs in future information retrieval tasks, but the effects on adversarial documents remain unknown.
 On the other hand, specialized adversarial defenses, such as keyword filtering and decoding aggregation Xiang et al.
@@
 This trade-off highlights the need for a more nuanced defense mechanism that can distinguish between adversarial and benign documents without compromising retrieval quality.
-To address this challenge, we propose Graph-based Reranking against Adversarial Document Attacks (GRADA), a novel and effective defense framework designed to protect RAG systems from adversarial retrieval manipulations.
+To address this challenge, we propose Graph-based Reranking against Adversarial Document Attacks (GRADA), an effective defense framework designed to protect RAG systems from adversarial retrieval manipulations.
 Our key insight is that adversarial documents, while optimized for high query similarity, exhibit weaker semantic coherence with genuinely relevant documents in the retrieval set.
@@
 By propagating ranking scores through this graph structure, our approach prioritizes clusters of semantically consistent documents while suppressing adversarially crafted outliers.
-As illustrated in Figure 2, our method significantly enhances the robustness of RAG-based LLMs, mitigating adversarial influences while preserving the integrity of benign retrieval results.
-We conducted comprehensive experiments on Natural Questions (NQ), MS-MARCO, and HotpotQA across five different models.
+As illustrated in Figure˜2, our method significantly enhances the robustness of RAG-based LLMs, mitigating adversarial influences while preserving the integrity of benign retrieval results.
+We conducted comprehensive experiments on Natural Questions (NQ), MS-MARCO, and HotpotQA across six different models.
 Our method has shown at least a 30% decrease in reducing the Attack Success Rate (ASR), with improvements of up to 80% across various adversarial attack strategies.
 We summarize our contributions as follows:
-  • We introduce GRADA, a weighted similarity graph among retrieved documents iteratively propagates scores to mitigate the impacts of adversarial passages.
-  • We introduce a novel scoring function that simultaneously captures both query-document and document-document correlations, thereby improving robustness against adversarial attempts to mimic the query.
-  • Conducted comprehensive experiments across three different datasets against four chosen attacks. Showing GRADA’s advantages over the current defense baselines.
+  • We introduce GRADA, which constructs a weighted similarity graph among retrieved documents and iteratively propagates scores to mitigate the impacts of adversarial passages.
+  • We introduce a novel scoring function that simultaneously considers both query-document and document-document correlations, thereby improving robustness against adversarial attempts to mimic the query.
+  • We conducted comprehensive experiments on three distinct datasets, evaluating our method against four representative attack types.
+The results consistently demonstrate that GRADA outperforms existing defense baselines.
 2 Related Work
-The IR community has long examined how malicious actors distort ranking functions.
-Gyongyi and Garcia-Molina (2005) propose a taxonomy of web-spam strategies, categorizing attacks into content-based, link-based, and behavior-based manipulations.
-Content-based spam involves manipulating the textual content of web pages to boost rankings, such as stuffing keywords or inserting irrelevant but popular terms.
-Link-based spam, including techniques like link farms, exploits the graph structure of the web to artificially inflate a page’s authority.
-Ntoulas et al.
-(2006) focus on detecting such content-based spam using statistical features extracted from term distributions and document structure.
-Their work highlights how spam pages often exhibit unnatural language models and disproportionately high keyword densities.
-Castillo and Davison (2011) provide a comprehensive survey of adversarial techniques in web search, including cloaking (serving different content to users and crawlers), redirection, and deceptive anchor texts.
-These traditional adversarial IR strategies, although originally developed to mislead search engines, reveal core vulnerabilities in retrieval systems—vulnerabilities that persist, albeit in more complex forms, in modern neural IR and RAG systems.
+Adversarial manipulation in IR has a long history.
+Gyongyi and Garcia-Molina (2005) categorize web-spam strategies into content-based, link-based, and behavior-based attacks, while Ntoulas et al.
+(2006) use statistical features to detect spam content.
+Castillo and Davison (2011) survey a range of traditional attacks like cloaking and redirection, which expose fundamental weaknesses that persist in modern neural retrieval systems.
 When RAG systems came out, Corpus poisoning attacks Zhong et al.
-(2023) show a possible new attack surface on LLMs.
-However, this method does not directly affect the accuracy of the LLM; instead, it focuses on the retriever.
+(2023) and third-party API attacks Zhao et al.
+(2024) show a new potential attack surface on LLMs.
 Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al.
@@
 (2024).
-However, compared to the prior work, these methods are unstable in ensuring the retriever retrieves the adversarial passage every time.
-While these attacks are situated in modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
+However, compared to the prior work, these methods are unstable in the retrieved adversarial passages.
+While these attacks are based on modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
 More recently, PoisonedRAG Zou et al.
 (2024) was proposed as a more stable attack.
-It uses two passages concatenated together, with one of them appended to guarantee the retrieval of the adversarial passage and one to achieve a given adversarial goal on the generator.
-The goal is to let the LLM output the answer the attacker wants.
-PoisonedRAG inspired a lot of the new attacks.
+It uses two passages concatenated together, with one of them appended to guarantee the retrieval of the adversarial passage and one to achieve the adversarial goal on the generator, which is to steer the LLM generating the answers anticipated by the attacker.
+PoisonedRAG inspired many subsequent attacks.
 Phantom Chaudhari et al.
 (2024), which introduces a trigger to the question and achieves the adversarial goal only when the trigger is shown in the query.
-Another Prompt Injection Attack (PIA, Clop and Teglia, 2024) makes use of the passage that guarantees the retrieval in PoisonedRAG and focuses on the adversarial goal beyond misinformation.
+Another type of Prompt Injection Attack (PIA, Clop and Teglia, 2024) leverages the guaranteed retrieval mechanism in PoisonedRAG.
+Unlike typical misinformation attacks, this variant targets broader adversarial objectives beyond merely spreading false information.
 A recent study proposed a defense mechanism that generates responses independently and produces an output based on the majority vote Xiang et al.
 (2024).
-However, this method initiates its defense at the generator stage, which can impact the accuracy of the system, especially when multiple documents are required.
-GRADA addresses this issue by focusing on the stage before generation, specifically the reranking process.
-Figure 3: BM25 similarity matrix among retrieved documents, where D0-D4 are poisoned, and D5-D10 are clean.
+While effective in some settings, this strategy defends only at the generator stage, which can compromise accuracy when multiple documents must be integrated.
+In contrast, several recent works intervene earlier in the RAG pipeline.
+TrustRAG Zhou et al.
+(2025) employs clustering and LLM-based conflict resolution to filter poisoned documents before they influence the generator, substantially reducing attack success rates while preserving accuracy.
+Complementary to this, traceback approaches such as RAGForensics Zhang et al.
+(2025) focus on identifying and removing the poisoned texts within the knowledge base itself, ensuring that subsequent retrieval yields only benign passages.
+Building on these insights, GRADA strengthens defenses at the reranking stage, preventing malicious content from reaching the generator without sacrificing the benefits of multi-document retrieval.
+Figure 3: The similarity matrix among retrieved documents using BM25, where D0-D4 are poisoned and D5-D10 are clean documents.
 3 GRADA
@@
 (2024) is their focus on ensuring semantic similarity to the query while introducing anomalous similarity patterns among poisoned documents.
-Adversarial documents are very similar to the query but less to the legitimate documents, forming isolated patterns within the retrieval set, as illustratedFigures 2 and 3.
+Adversarial documents closely resemble the query while diverging significantly from the legitimate documents, resulting in isolated patterns within the retrieval set, as illustrated in Figures˜2 and 3.
 Graph structures naturally capture these complex inter-document relationships by representing documents as nodes and similarities as edges.
 Leveraging this intuition, we propose a graph-based reranking method that utilizes document-document similarity to enhance retrieval robustness.
-In Section 3.1, we detail the graph construction process, followed by a description of our reranking system in Section 3.2.
+In Section˜3.1, we detail the graph construction process, followed by a description of our reranking system in Section˜3.2.
 3.1 Graph Construction
-We construct a weighted, undirected graph $G=(V,E)$, where each node $v_{i}\in V$ corresponds to a document $doc_{i}$, and each edge $e_{ij}\in E$ is an undirected edge connecting node $v_{i}$ and $v_{j}$.
-Each edge is assigned a weight $w_{ij}\in\mathbb{R}^{+}$, which quantifies the similarity between the corresponding documents, i.e., $\text{sim}(v_{i},v_{j})$.
+We construct a weighted, undirected graph $\mathcal{G}=(V,E)$, where each node $v_{i}\in V$ corresponds to a document, and each edge $e_{ij}\in E$ is an undirected edge connecting node $v_{i}$ and $v_{j}$.
+Each edge is assigned a weight $w_{ij}\in[0,1]$, which quantifies the similarity between the corresponding documents, i.e., $\text{sim}(v_{i},v_{j})$.
 The graph is undirected because document relationships are not inherently directional; rather, the connectivity structure defines their associations.
-The edge weight $w_{ij}$ can be computed using two different approaches:
+The edge weight $w_{ij}$ can be computed using different approaches as follows:
   • Doc-to-Doc Similarity (D2DSIM): The weight is directly determined by the similarity between documents.
   • Hybrid Relevance Similarity (HRSIM): A function $f$ that integrates both document-document similarity and query-document relevance:
-    $w_{ij}=f\big{(}\text{sim}(v_{i},v_{j}),\text{sim}(v_{i},q),\text{sim}(v_{j},q)\big{)}$
+    $w_{ij}=f\big{(}\text{sim}(v_{i},v_{j}),\text{sim}(v_{i},q),\text{sim}(v_{j},q)\big{)}.$
 The second approach assigns edge weights that not only reflect direct document-to-document similarity but also incorporate each document’s relevance to an external query.
@@
 First, we define the combined query relevance for a pair of documents $v_{i}$ and $v_{j}$ as follows:
-$\text{sim}_{\text{sum}}=\text{sim}(v_{i},q)+\text{sim}(v_{j},q)$
+$\text{sim}_{\text{sum}}=\text{sim}(v_{i},q)+\text{sim}(v_{j},q).$
 Then, the edge weight $w_{ij}$ between $v_{i}$ and $v_{j}$ is computed by subtracting a penalty term from their direct similarity, ensuring that the weight remains non-negative:
-$w_{ij}=\max\left(\text{sim}(v_{i},v_{j})-\alpha\cdot\text{sim}_{\text{sum}},0\right)$
+$w_{ij}=\max\{\text{sim}(v_{i},v_{j})-\alpha\cdot\text{sim}_{\text{sum}},0\}.$
 Here, $\alpha$ is a penalty coefficient that controls the influence of query similarity.
-If $\text{sim}(v_{i},v_{j})<\alpha\cdot\left[\text{sim}\left(v_{i},q\right)+\text{sim}\left(v_{i},q\right)\right]$, the edge weight is set to zero, effectively removing the connection between $v_{i}$ and $v_{j}$.
+If $\text{sim}(v_{i},v_{j})<\alpha\cdot\left[\text{sim}\left(v_{i},q\right)+\text{sim}\left(v_{i},q\right)\right]$, the edge weight is set to zero, which indicates removing the collusion connection between $v_{i}$ and $v_{j}$.
 Regarding the similarity function, we explore two popular methods:
@@
 Since BM25 is an asymmetric metric, we adopt the following approach to compute the similarity score, ensuring symmetry in the process:
-    $w_{ij}=\frac{1}{2}\left(\text{BM25}\left(v_{i},v_{j}\right)+\text{BM25}\left(v_{j},v_{i}\right)\right)$
+    $w_{ij}=\frac{1}{2}\left(\text{BM25}\left(v_{i},v_{j}\right)+\text{BM25}\left(v_{j},v_{i}\right)\right).$
   • Embedding-based Distance (EBD): we transform the documents $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ into dense vectors $v_{i}$ and $v_{j}$ and compute their cosine distance:
-    $w_{ij}=\text{sim}(v_{i},v_{j})=\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|}$
+    $w_{ij}=\text{sim}(v_{i},v_{j})=\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|}.$
 3.2 Reranking
@@
 The scores are then iteratively updated at each step $t$ via:
-$s_{i}^{(t)}$ $=(1-d)s^{*}_{j}+d\sum_{v_{j}\in\mathcal{N}(i)}\frac{w_{ij}}{\sum_{v_{k}\in\mathcal{N}(j)}w_{jk}}s_{j}^{(t-1)}$ (1)
-where $N(i)$ represents the set of neighbor nodes connected by $v_{i}$ and $d$ is the damping factor, typically set to 0.85.
+$s_{i}^{(t)}$ $=(1-d)s^{*}_{j}+d\sum_{v_{j}\in\mathcal{N}(i)}\frac{w_{ij}}{\sum_{v_{k}\in\mathcal{N}(j)}w_{jk}}s_{j}^{(t-1)},$ (1)
+where $\mathcal{N}(i)$ represents the set of neighbor nodes connected by $v_{i}$ and $d$ is the damping factor, typically set to 0.85.111The experiments comparing different initialization methods are provided in Section C.3.
 The initial score vector $s^{*}$ is set by uniform initialization $s^{*}=\left[\frac{1}{|V|},\frac{1}{|V|},...,\frac{1}{|V|}\right]$.
-For experiments comparing different initialization methods, please refer to Section C.2.
-The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval and $M\geq n$.
-We retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
-By ensuring that poisoned documents do not form the majority in the retrieved set (with $M\geq 2n$), we prevent adversarial documents, which may exploit the query for high relevance scores, from dominating.
-For example, if the original set of $n$ documents contains all poisoned ones (e.g., $n=5$), adding $\geq n$ benign documents ensures the majority is non-poisoned.
-This strategy guarantees that non-poisoned documents remain a significant portion of the final selection, enhancing the system’s robustness against adversarial manipulation.
+The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval (usually $M\geq n$).
+In experiments, we retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
+To prevent adversarial documents from dominating the retrieved set, we ensure that poisoned documents do not constitute the majority by retrieving at least twice the number of documents (i.e., $M\geq 2n$).
+For instance, if all $n$ original documents are poisoned (e.g., $n=5$), incorporating at least $n$ additional benign documents guarantees that the majority of the final selection is non-poisoned.
+This approach maintains a substantial presence of benign content in the retrieved set, thereby improving the system’s resilience to adversarial manipulation.
 After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded.
-These top $n$ documents are then provided as the context of the model.
+Then, these top $n$ documents are provided as the context of the model.
 4 Experiments
@@
 (2024).
-The prompts used to generate answers are detailed in Appendix A.
+The prompts used to generate answers are detailed in Appendix˜A.
 Contriever Izacard et al.
@@
 Additionally, this process is repeated using 3 random seeds, meaning each attack-defense pair is evaluated on 900 questions in total.
-However, in contrast to Zou et al.
-(2024), where five poisoned texts are generated and injected into the knowledge base, To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database.
+In contrast to Zou et al.
+(2024), where five poisoned texts are generated and injected into the knowledge base.
+To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database.
 The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack.
-As shown in Figure 3, a similarity cluster of poisoned documents appears in the top-left corner.
+As shown in Figure˜3, a similarity cluster of poisoned documents appears in the top-left corner.
 By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries.
 This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
-Defense setup.
-We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function.
-[^1] We examine other similarity functions in Section 4.3 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024).
-We compare GRADA against two reranking models and one defense method: HLATR Zhang et al.
-(2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al.
-(2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al.
-(2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
-We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section 3.
-We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation).
-The top five ranked documents are subsequently provided as the context for the model to answer the query.
-This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question.
-For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
-Evaluation metrics.
-In our experiments, we employ Attack Success Rate (ASR) and Exact Match (EM) as metrics.
-ASR is defined as the ratio of successful attacks to the total number of attacks conducted.
-An attack is considered successful if the intended poisoned answer appears as a substring within the generated response from the model.
-This definition accommodates attack strategies like PIA, which aim to introduce harmful links into the output of the model, allowing for some tolerance to semantically equivalent responses.
-A higher ASR indicates a more successful attack.
-This evaluation methodology follows the approach used in previous work Zou et al.
-(2024).
-To assess the question-answering accuracy of the models, we adopt EM score.
-EM requires that the predicted answer of the model matches the ground truth answer exactly.
-This strict criterion ensures that the response of the model is precise and follows the need for exact wording specified in the query, as outlined in Appendix A.
 Defense PoisonedRAG PIA
@@
 Table 1: ASR and EM (%) for various defense methods on the black-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct.
-The results of other models can be found in Tables 10, 9, 11, 13 and 12.
+The results of other models can be found in Tables˜11, 10, 12, 14 and 13.
 We highlight the top-2 lowest ASR results in blue cells.
+Defense setup.
+We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function.
+[^2] We examine other similarity functions in Section C.2 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024).
+We compare GRADA against two reranking models and one defense method: HLATR Zhang et al.
+(2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al.
+(2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al.
+(2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
+We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section˜3.
+We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation).
+The top five ranked documents are subsequently provided as the context for the model to answer the query.
+This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question.
+For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
+Evaluation metrics.
+In our experiments, we employ Attack Success Rate (ASR) and Exact Match (EM) as metrics.
+ASR is defined as the ratio of successful attacks to the total number of attacks conducted.
+An attack is considered successful if the intended poisoned answer appears as a substring within the generated response from the model.
+This definition accommodates attack strategies like PIA, which aim to introduce harmful links into the output of the model, allowing for some tolerance to semantically equivalent responses.
+A higher ASR indicates a more successful attack.
+This evaluation methodology follows the approach used in previous work Zou et al.
+(2024).
+To assess the question-answering accuracy of the models, we adopt EM score.
+EM requires that the predicted answer of the model matches the ground truth answer exactly.
+This strict criterion ensures that the response of the model is precise and follows the need for exact wording specified in the query, as outlined in Appendix˜A.
 Defense PoisonedRAG(Hotflip) Phantom
@@
 Table 2: ASR and EM (%) for various defense methods on White-box attacks.
-4.2 Results and Discussions
-Attacking without defense.
-As shown in Table 1, including a single poisoned document in the retrieval process results in a high ASR score.
-For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct.
-PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo.
-These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
-Effectiveness of GRADA.
-The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables 1 and 2.
-As shown in Table 1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD.
-With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%.
-However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks.
-On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%.
-Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
-The defense methods demonstrate consistent effectiveness across the NQ and MS-MARCO datasets, achieving ASR reductions of over 30% in most cases.
-However, performance on HotpotQA is less stable, particularly for D2DSIM-EBD and D2DSIM-BM25, which achieve only around a 10% reduction in ASR against PoisonedRAG attacks.
-In contrast, HRSIM maintains its effectiveness, delivering ASR reductions exceeding 30%, comparable to its performance on other datasets.
-This discrepancy likely stems from HotpotQA’s multi-hop reasoning requirements, which pose challenges for single-document similarity metrics.
-In Table 1, HLATR and BGE-reranker exhibit limited ability to filter poisoned documents, with ASR remaining largely unchanged compared to scenarios without any defense mechanisms.
-Although BGE-reranker occasionally outperforms HLATR, its overall performance remains inferior to GRADA in handling adversarial cases.
-This discrepancy underscores a critical limitation in contemporary reranking systems, which are primarily optimized for question relevance but insufficiently equipped to address adversarial attacks with high question relevance.
-Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom.
-Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements.
-However, while it reduces ASR effectively, its EM scores are lower than those of GRADA.
-For example, on Llama3.1-8b-Instruct in Table 1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference as some critical information may be lost during keyword extraction.
-This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
-Similar results to those presented in Table 1 can also be observed in Table 2.
-Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
 Defense HotpotQA NQ MS-MARCO
@@
 Table 3: EM scores of defense methods when presented with benign inputs.
-Table 3 highlights the impact of different defense mechanisms on benign inputs.
+4.2 Results and Discussions
+Attacking without defense.
+As shown in Table˜1, including a single poisoned document in the retrieval process results in a high ASR score.
+For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct.
+PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo.
+These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
+Effectiveness of GRADA.
+The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables˜1 and 2.
+As shown in Table˜1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD.
+With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%.
+However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks.
+On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%.
+Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
+The defense methods demonstrate consistent effectiveness across the NQ and MS-MARCO datasets, achieving ASR reductions of over 30% in most cases.
+However, performance on HotpotQA is less stable, particularly for D2DSIM-EBD and D2DSIM-BM25, which achieve only around a 10% reduction in ASR against PoisonedRAG attacks.
+In contrast, HRSIM maintains its effectiveness, delivering ASR reductions exceeding 30%, comparable to its performance on other datasets.
+This discrepancy likely stems from HotpotQA’s multi-hop reasoning requirements, which pose challenges for single-document similarity metrics.
+In Table˜1, HLATR and BGE-reranker exhibit limited ability to filter poisoned documents, with ASR remaining largely unchanged compared to scenarios without any defense mechanisms.
+Although BGE-reranker occasionally outperforms HLATR, its overall performance remains inferior to GRADA in handling adversarial cases.
+This discrepancy underscores a critical limitation in contemporary reranking systems, which are primarily optimized for question relevance but insufficiently equipped to address adversarial attacks with high question relevance.
+Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom.
+Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements.
+However, while it reduces ASR effectively, its EM scores are lower than those of GRADA.
+For example, on Llama3.1-8b-Instruct in Table˜1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference, as some critical information may be lost during keyword extraction.
+This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
+Similar results to those presented in Table˜1 can also be observed in Table˜2.
+Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
+Table˜3 highlights the impact of different defense mechanisms on benign inputs.
 On GPT-3.5-Turbo, both HLATR and BGE-reranker demonstrate strong performance, outperforming GRADA and enhancing the model’s overall accuracy.
@@
 GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 4%.
-Notably, the use of D2DSIM-BM25 leads to an 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall.
-However, on HotpotQA, HRSIM resulted in an 14% reduction in EM scores when handling benign inputs.
+Notably, the use of D2DSIM-BM25 leads to a 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall.
+However, on HotpotQA, HRSIM resulted in a 14% reduction in EM scores when handling benign inputs.
 While this trade-off is significant, it corresponds to HRSIM’s remarkable ASR reduction.
 Striking a balance between retrieval quality and defense robustness remains a crucial challenge for future research.
-Figure 4: Distribution of poisoned document positions after applying GRADA (HRSIM) in the NQ dataset.
-Documents positioned below rank 5 are effectively mitigated by the ranking algorithm.
-Other results are showed in Figures 11, 6, 7 and 8
 Keyword Aggregation has a much lower performance also in EM scores on benign input compared to GRADA.
@@
 Why GRADA works.
-For the attack to be effective, the attackers must ensure that the retriever selects the poisoned documents.
-To achieve this, they primarily focus on making these documents resemble the queries, as most retrieval models prioritize query-document similarity when selecting relevant results.
-Additionally, poisoned documents typically exhibit only weak similarity to other documents in the corpus.
-This characteristic makes them less susceptible to detection by defense mechanisms that compare retrieved documents against one another.
-GRADA exploits this by constructing a document similarity graph where documents effectively "vote" for other documents they strongly resemble semantically.
-Benign documents, which share common relevant content, naturally have strong mutual similarities (e.g., averaging  0.82) and thus reinforce each other.
-In contrast, adversarial documents—designed to mislead—tend to be more isolated, receiving fewer "votes" due to their significantly lower similarities (average  0.35) with genuine documents.
-As a result, GRADA promotes densely interconnected benign documents while reducing the influence of sparse adversarial content.
-A running example is also shown in Figure 14
-Figure 5: HRSIM performance with different similarity functions selection on MSMARCO dataset.
-The figure illustrates the proportion of test instances in which poisoned documents remain among the top five retrieved results.
-4.3 Additional Studies
-Ranking distribution.
-We have demonstrated the effectiveness of our approach in enhancing defense performance.
-To gain a deeper understanding of its impact, we further analyze how our method systematically lowers the ranking of poisoned documents.
-As illustrated in Figure 4, the position distribution of poisoned documents within the retrieval set shifts significantly after applying GRADA with D2DSIM-BM25.
-Notably, over 70% of poisoned documents are relegated beyond the top five positions, substantially reducing their influence.
-These findings confirm that GRADA is both robust and effective in mitigating adversarial attacks.
-Selections of HRSIM.
-Thus far, our focus has primarily been on utilizing BM25 for HRSIM.
-In this section, we explore other similarity functions for HRSIM.
-As shown in Figure 5, we extend our analysis by incorporating SBERT Reimers and Gurevych (2019), alongside the three previously discussed methods, to better capture document-to-document similarity.
-Our results indicate that both EBD and SBERT exhibit strong overall performance against PIA and PoisonedRAG attacks.
-In contrast, BGE-Reranker struggles to effectively filter out poisoned documents, likely due to its primary training objective of computing query-to-document similarities rather than document-to-document relationships.
-HRSIM, when combined with BM25, effectively minimize the presence of poisoned documents, reducing them to just 14 out of 100 test instances.
-This outcome underscores its remarkable effectiveness in filtering malicious content.
+For an attack to be effective, adversaries must ensure that the retriever selects the poisoned documents.
+To accomplish this, they typically craft these documents to closely resemble the query, exploiting the fact that most retrieval models prioritize query-document similarity.
+However, these adversarial documents often exhibit only weak similarity to the rest of the corpus, a property that makes them less likely to be flagged by defense mechanisms based on inter-document similarity comparisons.
+GRADA leverages this insight by constructing a document similarity graph in which each document effectively "votes" for other documents with which it shares strong semantic similarity.
+Benign documents, which naturally cluster around shared content, tend to form densely connected subgraphs with high mutual similarity (e.g., averaging $0.82$), thereby reinforcing each other.
+In contrast, poisoned documents—engineered to deceive—are typically more isolated, receiving fewer "votes" due to their low average similarity to genuine documents (e.g., $0.35$).
+As a result, GRADA amplifies the collective influence of benign content while attenuating the impact of sparsely connected adversarial documents.
+A running example is provided in Figure˜13.
 Figure 4: Impact of the $\alpha$ value as it increases with three metrics (ASR, number of poisoned documents, and EM) on NQ dataset with GPT-3.5-Turbo.
 Impact of $\alpha$ and $M$.
-As shown in Figure 6, the number of poisoned documents in the context decreases as $\alpha$ increases, reaching a minimum at $\alpha=0.3$ before starting to rise again after $\alpha=0.8$.
+As shown in Figure˜4, the number of poisoned documents in the context decreases as $\alpha$ increases, reaching a minimum at $\alpha=0.3$ before starting to rise again after $\alpha=0.8$.
 The ASR follows a similar trend to the number of poisoned documents after $\alpha=0.3$.
@@
 Figure 5: Impact of the $M$ value as it changes with three metrics (ASR, number of poisoned documents, and EM) on MSMARCO dataset with GPT-3.5-Turbo.
-Defense Total Time (s) Processing Time (s) Defense-Only Time (s) Defense-Only Processing (s)
-Keyword Aggregation 12.61 11.11 11.59 9.21
-GRADA (D2DSIM-EBD) 1.56 1.12 0.62 0.62
-GRADA (D2DSIM-BM25) 0.97 0.52 0.02 0.02
-GRADA (HRSIM) 1.05 0.61 0.05 0.05
-Table 4: Runtime Comparison (on GPT-3.5-Turbo, average per query): Total Time (s) and Processing Time (s) represent the complete runtime for answering one question, including retrieval, defense method, and LLM response generation.
-In contrast, Defense-Only Time (s) measures exclusively the runtime of the defense methods themselves.
-Total Time is recorded using Python’s time.time() function, whereas Processing Time is measured with Python’s time.process_time() function.
-Figure 7 illustrates the effect of selecting $M=n$.
+Figure˜5 illustrates the effect of selecting $M=n$.
 It shows that, regardless of how documents are re-ranked, poisoned documents can still remain within the context provided to the model.
 However, this approach results in a 17% decrease in ASR and a 9% increase in EM, indicating that simply adjusting document positions can significantly impact model performance.
-This aligns with our observations in Table 5, and the specific positions of the documents are detailed in Figure 4.
+This aligns with our observations in Table˜4, and the specific positions of the documents are detailed in Figure˜10.
 By including additional documents for reranking and then retrieving only the top $n$ results, the ASR is further reduced from 21% to 10%, with only 14% of poisoned documents remaining in the context provided to the model.
 This demonstrates the importance of including extra documents during reranking to remove poisoned content and achieve better overall performance effectively.
-Computational Complexities.
-The overall complexity of GRADA consists of two main components:
-  • Similarity matrix construction: $O(N^{2})$, where N is the number of retrieved documents.
-This step can incur additional costs depending on the chosen similarity function.
-For example, using D2DSIM-EBD (embedding-based document similarity), the complexity becomes $O(N^{2}\cdot d)$, where $d$ is the embedding dimension.
-BM25-based similarity: the complexity is $O(N^{2}\cdot L)$, where L is the average document length.
-This is efficient due to the sparsity of token overlaps and inverted index optimizations.
-Here, since we are reranking the documents after the retrieval step.
-The retrieved documents set is usually constrained with limited amounts of data, making this a viable solution.
-  • Graph-based reranking (e.g., PageRank): $O(n+m)$, where n is the number of nodes (documents) and m is the number of edges in the constructed similarity graph.
-The only defense Keyword Aggregation requires querying the language model N times—once per document—to collect individual answers before aggregating: $O(N*C_{LM})$ (where $C_{LM}$ refers to the language model’s cost).
-This incurs significantly higher costs in terms of API calls and model generation time, especially with large models.
-GRADA, by comparison, does not require any model calls.
-The only required model call is after GRADA to query the final answer, making it more efficient and scalable for large-scale or production RAG deployments.
 5 Conclusion
@@
 Our study focuses on improving the robustness of RAG systems, thereby enhancing their reliability and minimizing harmful manipulations.
-We evaluated our proposed method, GRADA, using publicly available datasets as detailed in Appendix F.
+We evaluated our proposed method, GRADA, using publicly available datasets as detailed in Appendix˜F.
 We do not engage in harmful data practices.
-Acknowledgement
-Aryo Pradipta Gema was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics.
-Giwon Hong was supported by the ILCC PhD program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics.
-Xuanli He was funded by an industry grant from Cisco.
-Pasquale Minervini was partially funded by ELIAI, an industry grant from Cisco, and a donation from Accenture LLP.
-This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
 Appendix A Prompts to LLM
-Figure 8 shows the two prompts we use to query the LLMs.
+Figure˜6 shows the two prompts we use to query the LLMs.
 Prompts given to LLMs Prompt 1: You are a helpful assistant, below is a query from a user and some relevant contexts.
@@
 The Apple archive, housing important launch events, mentions a significant one on April 15, 2015.
-On this day, the compact, budget-friendly iPhone SE hit the market, impressing many gadget enthusiasts." Poisoned Document 2: "When did the apple iphone se come out.
+On this day, the compact, budget-friendly iPhone SE hit the market, impressing many gadget enthusiasts."
+Poisoned Document 2: "When did the apple iphone se come out.
 The iPhone series has released several models.
@@
 Despite its simplicity, PoisonedRAG degrades the accuracy of the LLMs significantly.
-As shown in Table 5 (first row), the attack achieves an ASR of 54.3% on average across three datasets with just one adversarial passage retrieved as the most similar to the query.
+As shown in Table˜4 (first row), the attack achieves an ASR of 54.3% on average across three datasets with just one adversarial passage retrieved as the most similar to the query.
 Attack Method HotpotQA NQ MS-MARCO Average
@@
 Our analysis reveals that the prepended query in the adversarial passage does not significantly affect the ASR.
-As shown in Table 5 (second row), removing the prepended query leads to an increase in the ASR.
+As shown in Table˜4 (second row), removing the prepended query leads to an increase in the ASR.
 This shows that the query was prepended only to ensure that the retriever retrieves the adversarial document, but not affecting the accuracy significantly.
-Furthermore, Table 5 (third and fourth row) shows that the position of the poisoned document within the retrieved documents set influences the ASR significantly, with a decrease in average ASR of 10%.
+Furthermore, Table˜4 (third and fourth row) shows that the position of the poisoned document within the retrieved documents set influences the ASR significantly, with a decrease in average ASR of 10%.
 This phenomenon is similar to the lost-in-the-middle effect Liu et al.
@@
 Due to its straightforward approach of prepending the query to the adversarial documents, PoisonedRAG attacks can be easily identified.
-As demonstrated in Figure 3 and Figure 9, the attacks injected into the database often exhibit considerable similarity to one another.
+As demonstrated in Figure˜3 and Figure˜7, the attacks injected into the database often exhibit considerable similarity to one another.
 By focusing on the similarities between the documents in the retrieved set, we can filter out adversarial passages and decrease the ASR.
@@
 Figure 8: Impact of increasing poisoned documents on GRADA’s performance in NQ dataset (GPT-3.5-Turbo, $M=10$).
-(a) D2DSIM-EBD
-(b) D2DSIM-BM25
-(c) HRSIM
-Figure 11: Distribution of Ground Truth document positions after applying GRADA in the NQ dataset with different ranking methods.
 Appendix C Ablation Study
 C.1 Number of poisoned documents increase
-As shown in Figure 10(a), the effectiveness of GRADA reduces as the proportion of poisoned documents increases.
+As shown in Figure˜8(a), the effectiveness of GRADA reduces as the proportion of poisoned documents increases.
 When using D2DSIM-EBD, the ASR achieved by GRADA approaches that of an undefended system.
 However, HRSIM remains effective, achieving a 27% reduction in ASR even when half of the retrieved documents are adversarial.
-This is further supported by Figure 10(b), which shows that 38% of poisoned documents are still successfully filtered.
-C.2 Different initial score vector
+This is further supported by Figure˜8(b), which shows that 38% of poisoned documents are still successfully filtered.
+Figure 9: HRSIM performance with different similarity functions selection on MSMARCO dataset.
+The figure illustrates the proportion of test instances in which poisoned documents remain among the top five retrieved results.
+Figure 10: Distribution of poisoned document positions after applying GRADA (HRSIM) in the NQ dataset.
+Documents positioned below rank 5 are effectively mitigated by the ranking algorithm.
+Other results are showed in Figures˜14, 7, 8 and 9
+C.2 Selections of HRSIM.
+Thus far, our focus has primarily been on utilizing BM25 for HRSIM.
+In this section, we explore other similarity functions for HRSIM.
+As shown in Figure˜9, we extend our analysis by incorporating SBERT Reimers and Gurevych (2019), alongside the three previously discussed methods, to better capture document-to-document similarity.
+Our results indicate that both EBD and SBERT exhibit strong overall performance against PIA and PoisonedRAG attacks.
+In contrast, BGE-Reranker struggles to effectively filter out poisoned documents, likely due to its primary training objective of computing query-to-document similarities rather than document-to-document relationships.
+HRSIM, when combined with BM25, effectively minimize the presence of poisoned documents, reducing them to just 14 out of 100 test instances.
+This outcome underscores its remarkable effectiveness in filtering malicious content.
+C.3 Different initial score vector
 Different initial score vectors can have a significant impact on the final distribution of documents in certain cases.
 For instance, we experimented with initializing the score vector with query-document similarity $s^{*}=\left[\frac{sim(q,v_{0})}{\sum_{j=0}^{n}sim(q,v_{j})},\frac{sim(q,v_{1})}{\sum_{j=0}^{n}sim(q,v_{j})},...,\frac{sim(q,v_{n})}{\sum_{j=0}^{n}sim(q,v_{j})}\right]$.
-As shown in Figure 12(a), using a query-document initialization results in more documents being positioned between rank 5 and 8, rather than lower.
+As shown in Figure˜11(a), using a query-document initialization results in more documents being positioned between rank 5 and 8, rather than lower.
 We hypothesize that this is because adversarial documents may receive disproportionately high initial scores compared to benign documents.
 Such an imbalance gives adversarial documents a substantial advantage, particularly when the edge weights between documents are relatively small.
-In these scenarios, the graph-based reranking process may struggle to compensate for this initial disparity, as illustrated in Figure 14.
-From the analysis in Figure 12(b), we observe that this phenomenon is more prevalent in datasets like HotpotQA.
+In these scenarios, the graph-based reranking process may struggle to compensate for this initial disparity, as illustrated in Figure˜13.
+From the analysis in Figure˜11(b), we observe that this phenomenon is more prevalent in datasets like HotpotQA.
 (a) Distribution of Poisoned document positions after applying GRADA (HRSIM) with different initialization in the NQ dataset.
@@
 Figure 11: Impact of different initialization score vectors on GRADA’s performance ($M=10$).
+C.4 Ranking distribution.
+We have demonstrated the effectiveness of our approach in enhancing defense performance.
+To gain a deeper understanding of its impact, we further analyze how our method systematically lowers the ranking of poisoned documents.
+As illustrated in Figure˜10, the position distribution of poisoned documents within the retrieval set shifts significantly after applying GRADA with D2DSIM-BM25.
+Notably, over 70% of poisoned documents are relegated beyond the top five positions, substantially reducing their influence.
+These findings confirm that GRADA is both robust and effective in mitigating adversarial attacks.
+Defense Total Time (s) Processing Time (s) Defense-Only Time (s) Defense-Only Processing (s)
+Keyword Aggregation 12.61 11.11 11.59 9.21
+GRADA (D2DSIM-EBD) 1.56 1.12 0.62 0.62
+GRADA (D2DSIM-BM25) 0.97 0.52 0.02 0.02
+GRADA (HRSIM) 1.05 0.61 0.05 0.05
+Table 5: Runtime Comparison (on GPT-3.5-Turbo, average per query): Total Time (s) and Processing Time (s) represent the complete runtime for answering one question, including retrieval, defense method, and LLM response generation.
+In contrast, Defense-Only Time (s) measures exclusively the runtime of the defense methods themselves.
+Total Time is recorded using Python’s time.time() function, whereas Processing Time is measured with Python’s time.process_time() function.
+C.5 Different Retriever
+To further assess the generality of our approach, we conduct experiments on bge-small-en-v1.5 Xiao et al.
+(2023).
+As shown in Table˜6, the overall trends remain consistent with our main experiments.
+In the no-attack setting, all defenses incur only a small EM drop compared to the baseline retriever, suggesting limited utility loss.
+Under PoisonedRAG, the baseline suffers an ASR of 41.3% with a large EM drop, while our defenses substantially reduce ASR and recover EM—HRSIM achieves the best trade-off with an ASR of only 5.3% and the highest EM of 62.3%.
+For PIA, the baseline reaches 31.0% ASR, but our methods again nearly eliminate the attack, with D2DSIM-EBD completely blocking it while retaining 59.0% EM.
+These results demonstrate that our defenses transfer effectively across retrievers and maintain robustness against different poisoning strategies.
+Defense No Attack PoisonedRAG PIA
+ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$ ASR $\downarrow$ / EM $\uparrow$
+[…]
+[…]
+[…]
+[…]
+Table 6: Performance of defenses on an bge-small-en-v1.5 under No Attack, PoisonedRAG, and PIA.
+C.6 Computational Complexities.
+The overall complexity of GRADA consists of two main components:
+  • Similarity matrix construction: $O(N^{2})$, where N is the number of retrieved documents.
+This step can incur additional costs depending on the chosen similarity function.
+For example, using D2DSIM-EBD (embedding-based document similarity), the complexity becomes $O(N^{2}\cdot d)$, where $d$ is the embedding dimension.
+BM25-based similarity: the complexity is $O(N^{2}\cdot L)$, where L is the average document length.
+This is efficient due to the sparsity of token overlaps and inverted index optimizations.
+Here, since we are reranking the documents after the retrieval step.
+The retrieved documents set is usually constrained with limited amounts of data, making this a viable solution.
+  • Graph-based reranking (e.g., PageRank): $O(n+m)$, where n is the number of nodes (documents) and m is the number of edges in the constructed similarity graph.
+The only defense Keyword Aggregation requires querying the language model N times—once per document—to collect individual answers before aggregating: $O(N*C_{LM})$ (where $C_{LM}$ refers to the language model’s cost).
+This incurs significantly higher costs in terms of API calls and model generation time, especially with large models.
+GRADA, by comparison, does not require any model calls.
+The only required model call is after GRADA to query the final answer, making it more efficient and scalable for large-scale or production RAG deployments.
 Appendix D Different initial score vector demonstration
-Initial Score Example in Figure 14 Question: "Are Random House Tower and 888 7th Avenue both used for real estate?" Documents 1: "The former Bertelsmann Building, now known as 1540 Broadway, is a 44-story, 733 foot (223 m) office tower in Times Square in Manhattan…" Documents 2: "The Random House Tower, also known as the Park Imperial Apartments, is a 52-story mixed-use tower in New York City, United States, that is…" Documents 3: "888 7th Avenue is a 628 ft (191m) tall modern-style office skyscraper in Midtown Manhattan which was completed in 1969 and has 46 floors.
+Initial Score Example in Figure˜13 Question: "Are Random House Tower and 888 7th Avenue both used for real estate?" Documents 1: "The former Bertelsmann Building, now known as 1540 Broadway, is a 44-story, 733 foot (223 m) office tower in Times Square in Manhattan…" Documents 2: "The Random House Tower, also known as the Park Imperial Apartments, is a 52-story mixed-use tower in New York City, United States, that is…" Documents 3: "888 7th Avenue is a 628 ft (191m) tall modern-style office skyscraper in Midtown Manhattan which was completed in 1969 and has 46 floors.
 Emery Roth & Sons designed…" Documents 4: "What do the estates of film stars Vincent Price and Glenn Ford have in common?
@@
 888 7th Avenue is primarily used for law firms, again not real estate operations."
-Figure 13: Document examples used to generate Figure 14 to demonstrate different initial score vector and their results when the adversarial documents receive significantly higher initial scores compared to benign documents.
+Figure 12: Document examples used to generate Figure˜13 to demonstrate different initial score vector and their results when the adversarial documents receive significantly higher initial scores compared to benign documents.
 Red Documents indicates the poisoned document.
-Figure 13 shows the documents we used in Figure 14.
+Figure˜12 shows the documents we used in Figure˜13.
 Figure 13: A demonstration on different initial score vector and their results when the adversarial documents receive significantly higher initial scores compared to benign documents.
@@
 The dataset used in our experiments is publicly available under Creative Commons Attribution 4.0 International (MS-MARCO) and Apache License 2.0 (NQ, HotpotQA).
-The code used in our experiments is publicly available under MIT License (BM25s, PoisonedRAG) https://github.com/agrzheng/GRADA
+The code used in our experiments is publicly available under MIT License (BM25s, PoisonedRAG) https://anonymous.4open.science/r/GRADA-266D
+(a) D2DSIM-EBD
+(b) D2DSIM-BM25
+(c) HRSIM
+Figure 14: Distribution of Ground Truth document positions after applying GRADA in the NQ dataset with different ranking methods.
 Defense Method PoisonedRAG PoisonedRAG(Hotflip) PIA Phantom
@@
 GRADA (HRSIM) 30.0±0.0 44.7±1.2 48.7±0.9
+Qwen2.5-14b-Instruct No-RAG 17.3±0.5 17.3±0.5 9.3±0.5
+None 45.7±0.5 45.7±0.5 45.3±0.5
+HLATR 34.7±0.5 34.7±0.5 34.3±0.5
+BGE-reranker 36.7±1.2 36.7±1.2 37.3±0.5
+Keyword Aggregation 17.3±0.5 17.3±0.5 9.7±0.5
+GRADA (D2DSIM-EBD) 33.0±0.8 33.0±0.8 38.7±0.5
+GRADA (D2DSIM-BM25) 43.7±0.5 43.7±0.5 36.3±0.9
+GRADA (HRSIM) 41.7±0.9 41.7±0.9 39.3±0.5
 Table 10: EM scores of defense methods on GPT-4o, Llama3.1-70b-Instruct and Qwen2.5-7b-Instruct when presented with benign inputs.
@@
 GRADA (HRSIM) 7.7±0.5 / 34.0±0.8 5.3±0.5 / 41.0±0.0 12.3±0.5 / 39.0±0.8
+Qwen2.5-14b-Instruct None 47.7±0.5 / 17.0±0.0 43.3±0.5 / 12.3±0.5 38.0±0.0 / 19.7±0.5
+HLATR 42.7±0.5 / 17.7±0.5 36.3±0.5 / 19.7±1.2 26.7±0.5 / 23.3±1.2
+BGE-reranker 43.0±0.0 / 19.7±0.5 35.3±0.5 / 23.7±0.5 25.0±0.0 / 22.7±0.5
+Keyword Aggregation 4.0±0.0 / 22.3±0.5 5.0±0.0 / 18.3±0.5 3.0±0.0 / 9.7±0.5
+GRADA (D2DSIM-EBD) 34.0±0.0 / 16.0±0.0 13.0±0.0 / 27.7±1.2 26.0±0.0 / 27.0±0.0
+GRADA (D2DSIM-BM25) 28.0±0.0 / 20.0±1.4 5.0±0.0 / 38.3±0.5 13.3±0.5 / 33.7±0.9
+GRADA (HRSIM) 7.3±0.5 / 21.3±0.9 1.0±0.0 / 40.0±0.0 8.3±0.5 / 36.3±1.2
 Table 11: ASR and EM (%) for various defense methods on PoisonedRAG on GPT-4o, Llama3.1-70b-Instruct and Qwen2.5-7b-Instruct. Blue cells indicate top-two lowest ASR.
@@
 GRADA (HRSIM) 6.0±0.0 / 33.0±0.0 4.3±0.9 / 45.3±0.5 10.7±0.5 / 39.0±1.4
+Qwen2.5-14b-Instruct None 50.0±1.6 / 20.0±0.8 47.3±2.4 / 16.0±0.8 42.3±0.9 / 20.0±2.2
+HLATR 47.3±1.2 / 17.3±1.2 37.0±0.8 / 21.0±2.2 28.3±1.2 / 28.0±1.6
+BGE-reranker 43.3±2.9 / 23.7±1.7 32.7±0.5 / 27.7±2.9 27.3±0.5 / 25.7±1.2
+Keyword Aggregation 4.3±0.5 / 22.3±0.9 5.0±0.0 / 16.7±0.5 3.0±0.0 / 10.0±0.0
+GRADA (D2DSIM-EBD) 34.0±2.9 / 17.3±0.5 10.3±0.5 / 28.0±1.6 15.0±0.8 / 34.3±2.1
+GRADA (D2DSIM-BM25) 31.3±1.2 / 22.7±0.9 4.3±1.2 / 39.7±1.7 12.0±0.8 / 36.0±2.2
+GRADA (HRSIM) 9.0±0.0 / 22.0±0.0 1.3±0.5 / 42.3±0.5 7.7±0.9 / 37.0±0.8
 Table 12: ASR and EM (%) for various defense methods on PoisonedRAG(Hotflip). Blue cells indicate top-two lowest ASR.
@@
 GRADA (HRSIM) 8.7±0.5 / 35.3±0.5 2.0±0.0 / 46.3±0.9 1.0±0.0 / 47.3±1.7
+Qwen2.5-14b-Instruct None 99.0±0.0 / 0.0±0.0 94.0±0.0 / 3.0±0.0 87.0±0.0 / 6.7±0.5
+HLATR 98.0±0.0 / 0.0±0.0 88.7±0.5 / 3.0±0.0 83.0±0.0 / 5.7±0.5
+BGE-reranker 98.0±0.0 / 1.3±0.5 42.0±0.0 / 21.3±0.5 43.0±0.0 / 21.3±0.5
+Keyword Aggregation 0.0±0.0 / 23.0±0.0 0.0±0.0 / 18.7±0.5 0.0±0.0 / 9.7±0.5
+GRADA (D2DSIM-EBD) 33.0±0.0 / 14.0±0.0 2.0±0.0 / 29.0±1.4 3.0±0.0 / 37.3±0.5
+GRADA (D2DSIM-BM25) 40.7±0.5 / 17.0±0.0 12.0±0.0 / 35.3±0.5 2.0±0.0 / 37.3±0.5
+GRADA (HRSIM) 27.0±0.0 / 18.0±0.0 2.0±0.0 / 37.7±0.9 1.0±0.0 / 40.3±0.5
 Table 13: ASR and EM (%) for various defense methods on PIA. Blue cells indicate top-two lowest ASR.
@@
 GRADA (HRSIM) 14.0±2.4 / 27.3±0.9 0.7±0.5 / 36.3±0.5 0.0±0.0 / 48.7±1.7
+Qwen2.5-14b-Instruct None 67.7±2.1 / 0.3±0.5 51.0±1.4 / 4.0±1.6 43.7±2.5 / 16.3±2.1
+HLATR 72.0±2.9 / 1.0±0.0 56.0±2.2 / 3.7±0.9 52.3±3.4 / 12.7±0.5
+BGE-reranker 81.7±2.9 / 1.0±0.8 58.7±1.2 / 11.3±1.2 51.0±2.2 / 17.0±0.8
+Keyword Aggregation 0.0±0.0 / 12.7±0.5 0.0±0.0 / 11.3±1.2 0.0±0.0 / 7.0±0.0
+GRADA (D2DSIM-EBD) 44.7±3.1 / 5.0±0.8 12.0±2.2 / 24.7±2.5 11.0±2.2 / 34.7±0.5
+GRADA (D2DSIM-BM25) 23.7±0.9 / 18.7±0.5 5.0±0.8 / 34.0±1.4 0.3±0.5 / 38.0±0.8
+GRADA (HRSIM) 7.7±1.2 / 16.3±0.5 0.0±0.0 / 34.7±0.9 0.0±0.0 / 42.3±0.5
 Table 14: ASR and EM (%) for various defense methods on Phantom. Blue cells indicate top-two lowest ASR.

### v3

@@
 GRADA: Graph-based Reranking against Adversarial Documents Attack
-Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Author., Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
+Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Authors.
+Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, z.xuanli.he@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
 1 Introduction
@@
 However, compared to the prior work, these methods are unstable in the retrieved adversarial passages.
-While these attacks are based on modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
+While these attacks are based on modern LLM‑based retrieval, adversarial manipulation of information‑retrieval systems has a much longer history that is instructive for our setting.
 More recently, PoisonedRAG Zou et al.
@@
 The initial score vector $s^{*}$ is set by uniform initialization $s^{*}=\left[\frac{1}{|V|},\frac{1}{|V|},...,\frac{1}{|V|}\right]$.
-The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval (usually $M\geq n$).
-In experiments, we retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
+The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval.
 To prevent adversarial documents from dominating the retrieved set, we ensure that poisoned documents do not constitute the majority by retrieving at least twice the number of documents (i.e., $M\geq 2n$).
@@
 After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded.
-Then, these top $n$ documents are provided as the context of the model.
+Then, these top $n$ documents are provided as the context to the generative model.
 4 Experiments
-This section begins by detailing the experimental setup, followed by a comparison of our approach with multiple baseline methods.
-Finally, we analyze the effectiveness of our approach across different settings.
+This section begins with the experimental setup, followed by a comparison of our approach with multiple baseline methods. Finally, we compare and analyze our approach across different settings.
 4.1 Experimental Setup
@@
 Contriever Izacard et al.
-(2021), is a dense retriever model used to find relevant documents by calculating similarity scores between the query and the knowledge base.
-It was selected for this study due to its efficiency and ability to handle large datasets.
+(2021) is a dense retriever model used to find relevant documents by calculating similarity scores between the query and documents in the knowledge base.
+It was selected due to its efficiency and ability to handle large datasets.
 In this work, we investigate four distinct attack strategies on RAG.
@@
 (2024); Perez and Ribeiro (2022).
-The remaining two are white-box attacks, in which the attacker has access to the victim’s retriever: PoisonedRAG(Hotflip) Zou et al.
+The remaining two are white-box attacks, in which attackers have the access to the victim’s retriever, i.e., PoisonedRAG (Hotflip) Zou et al.
 (2024) and Phantom Chaudhari et al.
 (2024)
-Under default settings without any defense, as in Zou et al.
+Under default settings without defense, as in Zou et al.
 (2024), we retrieve the five most similar documents from the knowledge database to serve as the context for each question.
-We select 100 close-ended questions from each dataset, yielding 300 questions in total per attack-defense run.
+We select 100 close-ended questions from each dataset, yielding 300 questions in total per attack-defense experiment.
 Additionally, this process is repeated using 3 random seeds, meaning each attack-defense pair is evaluated on 900 questions in total.
-In contrast to Zou et al.
-(2024), where five poisoned texts are generated and injected into the knowledge base.
+In contrast, Zou et al.
+(2024) generated five poisoned texts and injected them into the retriever knowledge base.
 To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database.
 The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack.
-As shown in Figure˜3, a similarity cluster of poisoned documents appears in the top-left corner.
+As shown in Figure˜3, a similarity matrix appears to cluster poisoned documents in the top-left corner.
 By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries.
-This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
+This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack performance.
 Defense PoisonedRAG PIA
@@
 (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al.
-(2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
+(2024), the only existing defense specifically designed for RAG-based adversarial attacks.
 We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section˜3.
@@
 This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question.
-For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
+For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s keyword selections.
 Evaluation metrics.
@@
 GRADA (HRSIM) 8.7±1.7 / 44.7±1.2 2.0±0.8 / 53.3±2.6 6.3±0.9 / 72.3±1.2 10.3±2.1 / 41.3±1.9 0.3±0.5 / 53.7±0.9 0.0±0.0 / 60.3±1.7
-Table 2: ASR and EM (%) for various defense methods on White-box attacks.
+Table 2: Table 2: ASR and EM (%) for various defense methods on the white-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct.
 Defense HotpotQA NQ MS-MARCO
@@
 Attacking without defense.
-As shown in Table˜1, including a single poisoned document in the retrieval process results in a high ASR score.
-For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct.
+As shown in Table˜1, including a single poisoned document in the retrieval process results in a high ASR.
+For instance, PoisonedRAG achieves an ASR of around 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct.
 PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo.
@@
 As shown in Table˜1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD.
-With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%.
+With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing ASRs from 55.7% and 46.5% to 26.1% and 29.0%.
 However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks.
-On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%.
+In this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%.
 Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
@@
 Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements.
-However, while it reduces ASR effectively, its EM scores are lower than those of GRADA.
+While Keyword Aggregation reduces ASR effectively, its EM scores are usually lower than those of GRADA.
 For example, on Llama3.1-8b-Instruct in Table˜1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference, as some critical information may be lost during keyword extraction.
 This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
-Similar results to those presented in Table˜1 can also be observed in Table˜2.
+Similar results to those presented in Table˜1 can be observed in Table˜2 as well.
 Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.
@@
 These reranking systems yield at least a 2% improvement in EM scores, suggesting their effectiveness in mitigating noise unrelated to the posed questions.
-GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 4%.
+GRADA with D2DSIM-EBD effectively preserves model performance on benign inputs across all datasets, with EM score deviations remaining within 3%.
 Notably, the use of D2DSIM-BM25 leads to a 6% improvement in EM scores on NQ, matching the performance of BGE-reranker, which achieves the highest EM overall.
@@
 Keyword Aggregation has a much lower performance also in EM scores on benign input compared to GRADA.
-For example, in MS-MARCO, it results in 40% compared to 57% on Llama3.1-8b-Instruct and 59% compared to 75.3% on GPT-3.5-Turbo.
+For example, in MS-MARCO, it results in 36% compared to 57% on Llama3.1-8b-Instruct and 59% compared to 75.3% on GPT-3.5-Turbo.
 Indeed showing the cost of discarding valuable information when facing benign documents.
-Using GRADA, we demonstrate that it is possible to defend against the chosen attacks without compromising the model’s overall performance on EM.
+Using GRADA, we demonstrate that it is possible to defend against the chosen attacks without compromising the model’s overall performance.
 While reranking methods such as HLATR and BGE-reranker show promise in reducing noise, their limited effectiveness in countering adversarial attack noise highlights a critical gap in existing defenses.
-Similarly, Keyword Aggregation presents a valuable strategy for mitigating attack payloads but comes with trade-offs in EM scores.
+Similarly, Keyword Aggregation presents a valuable strategy for mitigating attack payloads but comes with significant trade-offs in EM scores.
 Why GRADA works.
-For an attack to be effective, adversaries must ensure that the retriever selects the poisoned documents.
+For effective attacks, adversaries should steer the retriever to select the poisoned documents.
 To accomplish this, they typically craft these documents to closely resemble the query, exploiting the fact that most retrieval models prioritize query-document similarity.
 However, these adversarial documents often exhibit only weak similarity to the rest of the corpus, a property that makes them less likely to be flagged by defense mechanisms based on inter-document similarity comparisons.
-GRADA leverages this insight by constructing a document similarity graph in which each document effectively "votes" for other documents with which it shares strong semantic similarity.
+GRADA leverages this insight by constructing a document similarity graph in which each document effectively “votes” for other documents with which it shares strong semantic similarity.
 Benign documents, which naturally cluster around shared content, tend to form densely connected subgraphs with high mutual similarity (e.g., averaging $0.82$), thereby reinforcing each other.
-In contrast, poisoned documents—engineered to deceive—are typically more isolated, receiving fewer "votes" due to their low average similarity to genuine documents (e.g., $0.35$).
+In contrast, poisoned documents—engineered to deceive—are typically more isolated, receiving fewer “votes” due to their low average similarity to genuine documents (e.g., $0.35$).
 As a result, GRADA amplifies the collective influence of benign content while attenuating the impact of sparsely connected adversarial documents.
-A running example is provided in Figure˜13.
-Figure 4: Impact of the $\alpha$ value as it increases with three metrics (ASR, number of poisoned documents, and EM) on NQ dataset with GPT-3.5-Turbo.
-Impact of $\alpha$ and $M$.
+A running example is provided in Figure˜13 in Appendix.
+Figure 4: Comparison of $\alpha$ on three metrics (ASR, number of poisoned documents, and EM), based on NQ dataset with GPT-3.5-Turbo.
+Impact of hyper-parameters $\alpha$ and $M$.
 As shown in Figure˜4, the number of poisoned documents in the context decreases as $\alpha$ increases, reaching a minimum at $\alpha=0.3$ before starting to rise again after $\alpha=0.8$.
@@
 When $\alpha=0.4$, all three metrics (ASR, number of poisoned documents, and EM) are within an acceptable range, approaching the optimal performance values for $\alpha$.
-Figure 5: Impact of the $M$ value as it changes with three metrics (ASR, number of poisoned documents, and EM) on MSMARCO dataset with GPT-3.5-Turbo.
+Figure 5: Comparison of $M$ value on three metrics (ASR, number of poisoned documents, and EM), based on MSMARCO dataset with GPT-3.5-Turbo.
 Figure˜5 illustrates the effect of selecting $M=n$.
@@
 5 Conclusion
-The study examines the robustness challenges faced by RAG systems.
+Our research examines the robustness challenges faced by RAG systems.
 We identify a critical vulnerability in current adversarial attacks, which focus on increasing semantic similarity to the query without accounting for the relationships between the retrieved documents.
@@
 We do not engage in harmful data practices.
+Acknowledgement
+Aryo Pradipta Gema was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics.
+Giwon Hong was supported by the ILCC PhD program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics.
+Xuanli He was funded by an industry grant from Cisco.
+Pasquale Minervini was partially funded by ELIAI, an industry grant from Cisco, and a donation from Accenture LLP.
+Qiongkai Xu acknowledges support from 2024 FSE Strategic Startup Grant.
+This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
 Appendix A Prompts to LLM